### v3

Abstract
[Abstract] Technical debt (TD) represents the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals. Self-Admitted Technical Debt (SATD) occurs when developers explicitly acknowledge these trade-offs in the codebase, typically through comments or annotations. Automated detection of SATD has become an increasingly important research area, particularly with the rise of natural language processing (NLP), machine learning (ML), and deep learning (DL) techniques that aim to streamline SATD detection.
This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and 2024, focusing on the evolution of techniques from NLP-based models to more advanced ML, DL, and Transformers-based models such as BERT. The review identifies key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1-score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and the explainability of models.
The findings suggest that while early NLP methods laid the foundation for SATD detection, more recent advancements in DL and Transformers models have significantly improved detection accuracy. However, challenges remain in scaling these models for broader industrial use. This SLR offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
[…]
keywords:
self-admitted technical debt, SATD, detection, identification, categorization, systematic review
articletype: Review Articlejournal: Journalvolume: 00
1 Introduction
Business needs often require developers to be able to deliver new products and features as rapidly as possible, in order to reach the market. In fact, a developer must be able to balance carefully designing applications and seamlessly coding their features. In the business or corporate world, however, schedules and limited resources frequently hinder developers from generating perfect code before deploying it into products 1. This becomes commonplace when stakeholders cut several technical processes to catch up on momentum, especially in startup practices 2, 3. A tight deadline frequently leads to technical debt, as developers lack time to consider long-term solutions 4.
Technical debt (TD) is often used to describe the trade-off that occurs when an organization (or a developer) chooses to prioritize the delivery of new features or functionality over the long-term maintenance of the software 5. TD can occur in any software development project and can be caused by a variety of factors, including lack of resources, time, knowledge, or experience on the part of the development team 6, 7.
Among various types of TD, self-admitted technical debt (SATD) has gained considerable attention due to its explicit acknowledgment within the source code comments by developers. SATD represents instances where developers intentionally or unintentionally leave behind suboptimal design or implementation choices, often with a clear understanding of the resultant negative impact on the codebase 8. This acknowledgment is not merely an admission of suboptimal practices but a vital piece of information that can guide future development activities. Recognizing SATD is crucial for maintaining software quality and sustainability, as it can accumulate over time and impede future development efforts.
Over the past decade, many approaches to detecting SATD have emerged, driven by advances in natural language processing (NLP), machine learning (ML), and deep learning (DL). These approaches aim to automate the detection of SATD, facilitate their management, and helping software teams address it proactively. Alongside these approaches, various methodologies have been proposed to detect SATD in different stages of the software development life cycle, leveraging manual and semi-automated, and automated techniques. However, the diversity of techniques, methodologies, tools, and evaluation metrics used in SATD detection has led to a fragmented landscape, making it challenging for researchers and practitioners to gain a holistic understanding of the field.
To address this fragmentation, we selected a study that aligns closely with our focus on SATD detection. We summarize this work, compare it with our systematic literature review (SLR), and provide a condensed analysis of the key differences. Our SLR distinguishes itself from prior reviews in several key areas:
Time frame and scope.
Our review is the most extensive to date, covering relevant studies published between 2014 and 2024. In contrast, the review by Sierra et al. focuses on a narrower time frame, covering studies published between 2014 and 2018 9.
Systematic approach.
We followed the original and rigorous procedures outlined by Kitchenham and Charters 10 for conducting systematic reviews. In contrast, Sierra et al.’s study 9 appears to be based on a survey approach, without explicitly detailing how studies were sourced.
Comprehensiveness.
In addition to relying on search engines, our review included a bidirectional snowballing process to ensure we captured all relevant articles. As a result, we analyzed a broader range of studies.
Profoundness.
Unlike Sierra et al. 9, who did not report the metric evaluation results, our review includes detailed analysis of metrics such as precision, recall, and F1-score, categorizing them based on the approaches and methodologies used.
Analysis.
While Sierra et al. 9 broadly discussed SATD activities without focusing deeply on detection approaches, our review provides a more detailed evaluation of detection techniques, their effectiveness, and the tools proposed for SATD detection.
In this study, we conduct and evaluate a systematic literature review that focuses on the detection of SATD, particularly using NLP-based approaches and their combinations. The objectives and key contributions of our study are as follows:
  • We consolidate and categorize various approaches employed to detect SATD, revealing various techniques and identifying critical trends over the past decade.
  • We examine methodologies proposed for SATD detection, showcasing their impact on automation and accuracy.
  • We investigate tools developed for SATD detection over the past decade, providing an overview of their effectiveness.
  • We compare the best-performing approaches by evaluating precision, recall, F1-score, and generalizability across different software systems and datasets.
  • We identify primary challenges in SATD detection, offering insights for future research and improvement.
In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section  3 describes the methodology utilized to conduct the SLR. Sections 4 and 5 summarize and discuss the obtained results, while Section 6 identifies the threats to validity of this study. Finally, Section 7 offers a conclusion.
2 Background
This section introduces the concepts underpinning this research, related to NLP and self-admitted technical debt, as well as the relevant works.
2.1 Self-Admitted Technical Debt (SATD)
The term “technical debt” was coined by software developer Ward Cunningham. He initially used the metaphor for communicating to non-technical stakeholders the importance of budgeting resources for refactoring 5. Technical debt, sometimes referred to as tech debt or code debt, is the result of a development team rushing the delivery of projects or features that must later be refactored. In other words, there is a trade-off between delivering high-quality, effective, and optimized code and the short-term benefits of achieving release dates 6, 4. A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 11, 12.
The term “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 13. It represents the debt incurred when development teams opt for quick and easy solutions in the short term, risking long-term negative impacts.
The term “TD” encompasses below-par aspects of the software that developers are aware of but lack resources to address, such as outdated documentation, unperformed tests, complex code needing refactoring, and known defects left unattended 14, 15. These immature artifacts can lead to challenges in meeting quality standards and cause unforeseen delays in implementing necessary changes. Numerous empirical studies have recently concentrated on examining source code comments and other textual artifacts that indicate technical debt, often known as self-admitted technical debt (SATD) 8. The source used to detect SATD is usually in the form of natural language often expressed through text, such as code comments, commit messages, or other textual artifacts; therefore, SATD detection significantly relies on NLP-based approaches.
2.2 Natural Language Processing (NLP)
Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language, encompassing tasks like speech recognition, natural language understanding, and natural language generation 16, 17. Chowdhury 17 explains that the objective of NLP researchers is to understand how people perceive and use language so that the appropriate tools and techniques can be developed to assist computers in understanding and using natural languages to complete the necessary tasks.
NLP deals with processing natural language, which is the language used by humans in text or audio form. Computational linguistics, rule-based human language modeling, statistical models, machine learning, and deep learning are combined to enable computer programs to interpret and process human language 18, 19. In the context of SATD detection, plain text is used as input, including project artifacts such as source code comments, commit messages, pull requests, and issue tracking systems 20, 21.
Source code comments serve as annotations within the code, providing documentation directly incorporated by programmers 22. These comments are essential for understanding a system and may contain important information, such as the purpose of code additions, progress status on collective tasks, or reasons for specific changes 23. Comments can express concerns that need attention in the future (e.g., the TODO keyword), identify emerging issues, and guide future actions 24, 25. They provide valuable context specific to their authors and are human-readable.
Recently, automated approaches utilizing NLP have gained traction in addressing software engineering problems, including the detection of TD, which is called SATD detection. NLP has been effectively used to automatically identify the presence of TD and even categorize specific types of SATD. Notably, numerous studies have integrated NLP with machine learning (ML) models or deep learning (DL) architectures and also Transformers, aiming to enhance model performance (evidence is provided in Subsection 4.1).
3 Research Methodology
This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al26, 27, 10. The procedures for SLR are broken down into three stages: planning, conducting, and reporting the SLR 10. We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure  1).
Figure 1: Graphical representation of the SLR process used in this article
The review process comprises several steps. Initially, research questions were formulated for the SLR. Subsequently, a search strategy was devised, involving the identification of search strings and the selection of online databases for locating primary studies. The third stage encompassed the identification of pertinent studies aligned with the research questions, along with the establishment of inclusion and exclusion criteria. Following this, bidirectional snowballing was conducted utilizing references from the source articles. After that, we created a data extraction form to answer the research questions. A data extraction form was then crafted to address the research questions. Lastly, a data synthesis approach was developed to analyze and summarize the research findings.
Figure 2 depicts the article selection procedure for determining the primary studies used in this SLR. The following subsections provide detailed information about the selection process.
Figure 2: An overview of the steps in determining primary studies
3.1 Research Questions
Our SLR includes all relevant research outlining the various SATD detection techniques. To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 28. Boolean operators were applied to structure the search effectively: OR was used to include synonyms and alternative terms, while AND linked the key terms, as shown in Table 1. This approach ensured a comprehensive and focused search strategy.
Table 1: PICOC Criteria
Parameter Description Keywords Used
Population Field of application “self-admitted technical debt” OR SATD
Intervention Approach or method for dealing with a certain problem “nlp” AND “technical debt”
Comparison Not aplicable Not aplicable
Outcome Related to factors of importance to practitioners “detection” OR “identification” OR “prediction”
Context Context in which intervention take places “software development” OR “software engineering”
According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect self-admitted technical debt. Therefore, the aim of this SLR is to answer the following questions:
RQ1: What approaches have been employed to detect self-admitted technical debt (SATD) in software development?
Rationale: This question aims to understand the progress made in detecting SATD. By surveying existing studies, we can detect trends and potential future research directions.
RQ2: What methodologies have been proposed to detect SATD in software development?
Rationale: This question aims to analyze the various approaches and tools developed over the past decade for SATD detection.
RQ3: How effective are these approaches regarding their metrics evaluation results?
Rationale: This question aims to evaluate the performance and practicality of the identified approaches by comparing their precision, recall, and F1-score.
RQ4: What tools have been proposed for SATD detection?
Rationale: This question aims to evaluate the current landscape of tools designed for detecting SATD from the last decade.
RQ5: What are the primary challenges associated with the detection of SATD?
Rationale: This question focuses on uncovering the common obstacles faced when detecting SATD, which could help direct future research and tool development.
3.2 Search Protocol
Before beginning the SLR planning stage, it is necessary to conduct a pilot search on a database such as Scopus to observe if there are any SLRs with the same research theme. There is no need to conduct a new SLR if there are existing SLRs on the same topic 29. In this systematic review of the literature, no results covering the topic of the capabilities of the NLP approach in identifying technical debt were discovered; therefore, the SLR can be conducted as new research.
The SLR’s search strategy consists of three stages: pilot search, primary search, and snowballing. In the pilot search, multiple search strings are used on online databases to determine the most effective one. The primary search uses the identified search string to find relevant literature in online source databases. The search concludes with a snowball search 30. The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore, ScienceDirect, ACM Digital Library, and Springer. These databases, as proved by previous software engineering SLRs 31, 32, 33, 34, provide an excellent overview of the software engineering literature and are trustworthy enough to cover our scope of the literature. Moreover, we employed the “Advanced Search” feature in these databases to obtain full-text and metadata results.
Table 2: General search strings utilized during pilot search
String ID Search String
S1 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”)
S2 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
S3 (“self-admitted technical debt" OR SATD) OR (“technical debt" AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering" OR “software development")
The results of a pilot search using search string S1 were communicated to the second author by the first author. The second author selected 5% articles at random to see if they were relevant. The second author proposed altering the search string to S2 because the search string S1 result was still too general. Similar to the first process, iteration is repeated until it is decided to utilize the final search string, namely search string S3.
As illustrated in Table 2, through a pilot search with multiple search strings, the most effective one (i.e., S3) was selected. The search string was designed with synonyms, alternative keywords, and boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies. Drawing from the pilot search conducted earlier, for this search string S3, we incorporated three supplementary terms commonly associated with tasks related to SATD identification: detect, identify, and predict. Additionally, wildcard asterisk (*) was used to broaden the search for variations of words 35.
The final search string (i.e., S3) was applied to the mentioned 6 online databases to find relevant primary studies for the research questions as described in Subsection 3.1. We utilized the “Advanced Search” feature in these databases to acquire full-text and metadata results.
Table 3: Search results from selected databases
No Database Search Result
1 Web of Science 51
2 Scopus 1,004
3 IEEE Xplore 168
4 ACM Digital Library 41
5 Springer 125
6 ScienceDirect 51
Total 1,440
As depicted in Figure 2, the initial step involved searching multiple online databases using predefined search queries to collect a comprehensive set of studies relevant to the research topic. As shown in Table 3), a total of 1,440 studies were retrieved at this stage. To manage such a large dataset, both authors worked collaboratively to remove duplicates, which resulted in 1,119 unique records. Following the duplicate removal, a filtering process based on metadata by using inclusion and exclusion criteria presented in Table 4 was conducted. At this stage, the titles, abstracts, and keywords of the remaining studies were examined to determine their relevance to SATD detection.
A full-text analysis of these 69 studies was then performed to evaluate the quality and relevance of the research in greater detail. The first author initially reviewed the metadata and conducted an independent assessment of the full papers, focusing on their contributions to SATD detection approaches. After completing the initial evaluation, the first author discussed the findings with the second author, who served as a supervisor throughout the process. This discussion provided a critical review of the selection criteria and the individual assessments.
To ensure that no bias influenced the selection process, the second author independently reviewed a subset of the full-text studies. These results were then compared with the first author’s evaluations, and any discrepancies or differences in judgment were carefully discussed and resolved. This collaborative approach minimized the risk of subjective bias and ensured that each study was rigorously and objectively evaluated based on its quality and relevance.
To further ensure consistency in the evaluation process, the test-retest strategy outlined by Kitchenham and Charters 10 was employed, rather than relying on inter-rater agreement. This strategy involved discussions between the first and second authors to maintain a consistent evaluation framework. Additionally, the second author cross-checked the data extraction results, which were shared in a spreadsheet [^1] The spreadsheet is available in the shared dataset area at https://github.com/edisutoyo/satd-detection-slr. Any disagreements that arose during the process were addressed through collaborative discussions until consensus was reached.
After this comprehensive review and resolution of all discrepancies, both authors agreed to select 63 relevant studies from 6 online databases to be included in the systematic review. This thorough, unbiased, and transparent process ensured the inclusion of only the most relevant and high-quality studies.
3.3 Article Selection and Snowballing Process
Inclusion and exclusion criteria must be precisely defined for study selection. The inclusion criteria should include an automated approach to detect technical debt. Consequently, the primary studies were chosen based on comparing titles, keywords, abstracts, and full texts (i.e., introduction, methodology, results, and conclusion). We employed 6 inclusion criteria and 5 exclusion criteria, as shown in Table 4):
Table 4: Inclusion and exclusion criteria
No Inclusion criteria
1 The article must be peer-reviewed and published at a conference, workshop, or journal
2 The article must be accessible online (i.e., articles can be fully accessed and downloaded)
3 For research that has both publication types (journal and proceeding version), only the journal version will be selected
4 The study must discuss approaches for identifying and measuring the performance of SATD detection
5 The study must be published between the years 2014 and 2024
6 The study reported must be related to SATD in software engineering or software development
No Exclusion Criteria
1 Articles written in languages other than English
2 Works in the form of dissertations, theses, and books
3 Articles that are still in the form of proposals
4 The article does not answer at least one research question
5 All other secondary and tertiary studies
To further enhance the comprehensiveness of the review, a bidirectional snowballing technique 30 was employed. Both forward and backward snowballing were performed, meaning that the references cited in the selected studies, as well as the studies that cited them, were examined to identify any additional relevant research that might have been overlooked in the initial search. This process resulted in the inclusion of 5 additional studies. The second author actively contributed to the snowballing process, assessing the relevance and quality of newly identified studies.
In the final stage, the 63 studies selected from the full-text filtering process were combined with the 5 studies identified through bidirectional snowballing, resulting in a total of 68 primary studies. This collaborative effort between the authors ensured that the selection process was both thorough and objective, ultimately leading to a comprehensive review of the existing literature on SATD detection approaches.
3.4 Quality Assessment Criteria
After finalizing the list of articles for the SLR, we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions. The quality of the selected articles was assessed using the questions presented in Table 5, following the procedure recommended by Kitchenham 26. All 68 relevant articles we discovered passed the quality evaluation, and thus, our SLR relies on these 68 primary studies.
Table 5: Quality assessment criteria
No Questions
1 Are the research objectives or research questions in the article clear?
2 Does the study include reviewing the relevant literature or background?
3 Does the paper describe a research methodology?
4 Does the article employ an algorithm or model for SATD detection?
5 Does the study provide research findings and answer the research questions?
6 Does the article report metric evaluation performance or at least report results of the approach?
7 Does the article clearly state the dataset used?
All of the listed articles were published between January 2014 and June 2024. Our searches for articles were completed by the end of June 2024, supplemented by articles obtained through bidirectional snowballing. As a result, it is highly unlikely that we missed any published articles during our period due to publication delays.
3.5 Data Extraction
We selected the final relevant articles after applying the inclusion and exclusion criteria and conducting the quality evaluation. We determined the required information and attributes for data summarization to address our research questions. This facilitated the data extraction process, where we read the full text of each article and extracted the relevant information based on specified attributes. To answer the research questions mentioned in Subsection 3.1, we extracted data items listed in Table 6 from each selected study. The extracted data was then entered into a spreadsheet for further analysis.
Table 6: Data extracted from each study
No Data Name Description Relevant RQ
1 Title Title of the study Overview
2 Author Name of the authors Overview
3 Year The publication year of the study Overview
4 Venue The name of the publication venue of the study Overview
5 Publication Type Journal, conference, workshop, or book chapter Overview
6 Algorithms used Algorithm or model used in the study RQ1, RQ2, RQ4
7 TD Types The types of TD, e.g., design or requirement TD R2, RQ3
8 Number of Class Number of class(es)/TD types used R2, RQ3
9 Dataset used Data source and metadata of dataset used R1, RQ4
10 Metrics Evaluation used Types of metric evaluations used for performance measurement RQ4
11 Performance Evaluations Performance measurement results using metric evaluations RQ4
12 Challenges Issues or challenges related to SATD detection RQ5
4 Results
This section addresses the posed research questions. To that purpose, the outcomes of the 68 primary studies are presented and discussed.
4.1 Results - RQ1: Approaches used for detecting SATD in software development
As part of the results for RQ1, we categorized the relevant studies into three main groups based on their approach: papers using NLP, those employing machine learning (ML), and studies utilizing deep learning (DL) or Transformers (see Figure 3 and Table 7). Any disagreements that arose during the grouping process were resolved through discussion.
Figure 3 shows the number of papers published by approach: NLP, ML, DL, and Transformers. Over the past decade, the identification of SATD has evolved significantly, driven by advancements in NLP and ML techniques. Early methods relied primarily on traditional NLP techniques, focusing on keyword detection and textual patterns in code comments. As depicted in Figure 3, NLP dominated initially, with limited exploration of advanced computational methods. From 2017 to 2019, ML approaches were increasingly introduced to improve SATD detection accuracy. Between 2020 and 2022, the use of DL and Transformer architectures surged, marking a shift toward more sophisticated, data-driven techniques. These models captured complex patterns in software artifacts, improving SATD detection. As of 2023 and 2024, DL and Transformer models continue to lead SATD detection efforts.
Figure 3: Articles count throughout the years based on their approaches
Table 7: Approach identified. Studies that fall into more than one approach are highlighted with square brackets
Approach Paper Reference Technique examples
NLP No. 8, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49 POS tagging, Regular Expressions, Textual patterns, etc.
Machine Learning No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64 65, 66, 67, 68, 69, 70, 71, 72,[73, 74, 75, 76] NB, ME, SVM, kNN, etc.
Deep Learning No. 77, 78, 79, 80, 81, 82, 82, 83, 84, 85, 86, 87, 88, 89 90, 91, 92, 93, 94,[73, 74, 75, 76] CNN, LSTM, BiLSTM, etc.
Transformers No. 95, 96, 97, 98, 99, 100,[73, 74, 75] BERT, ALBERT, GPT, etc.
NLP approaches have played a fundamental role in detecting SATD. These methods generally concentrate on text preprocessing tasks that encompass stemming, lemmatization, and stopwords removal. The objective is to process and evaluate textual data from code comments or documentation, where technical debt is frequently acknowledged by the authors themselves. As presented in Table 7, other methods include manual analysis and Parts-of-Speech (POS) tagging 36, 40, 41, 49, which aid in classifying and identifying relevant information. Regular expressions 38, 46, keyword-labeled 61, and textual patterns 45 are commonly employed to semi-automate the identification of certain technical debt indicators in large codebases.
As the complexity of software systems increased, so has the demand for more sophisticated approaches in technical debt identification, resulting in the use of ML classifiers. These algorithms include Naïve Bayes (NB) 50, 55, 59, Maximum Entropy (ME) 51, 67, Support Vector Machines (SVM) 52, and k-Nearest Neighbor (kNN). Logistic Regression (LR) 101 and Random Forest (RF) 56, 43 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt. Furthermore, XGBoost 66, 68, a relatively new addition to the ML classifier, improves performance by refining gradient boosting approaches.
The rise of DL has introduced even more advanced techniques for detecting SATD. Convolutional Neural Network (CNN) 77, 83, 85, 90, Long Short-Term Memory (LSTM) 79, 81, and Bidirectional Long Short-Term Memory (BiLSTM) 78, 82 are among the most prominent techniques used. They are particularly useful in processing structured data and identifying patterns that are not easily captured by traditional ML algorithms. Deep learning approaches have been extensively studied and are shown to be more critical in automatically detecting technical debt.
Transformers are the most recent development in the detection of SATD. By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 95, 96, 100, 97, A Lite BERT (ALBERT) 98, Robustly Optimized BERT (RoBERTa) 98, and DistilBERT 98, have transformed the field of NLP. Transformers excel at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt. The literature reflects a growing trend towards using these models to achieve higher accuracy and more robust identification processes, especially in dealing with large and complex datasets typical of modern software systems.
Answer to RQ1 Four main approaches have been identified to detecting SATD: traditional NLP, ML, DL, and Transformers. Over time, the focus has shifted from basic NLP techniques to more advanced methods like ML and DL, with transformers now leading the way, due to their effectiveness in capturing complex textual patterns.
4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
To address this, we categorized the methodologies into three main types: manual, semi-automated, and automated. These categories represent the evolution in improving SATD detection’s accuracy, efficiency, and scalability. Manual methods depend on human expertise to review text artifacts, while semi-automated techniques combine human input with tools to streamline the process. Fully automated methods use advanced algorithms to detect SATD with minimal human intervention. Each approach has its own strengths and challenges, which we will explore. Disagreements in categorization were resolved through discussion.
Figure 4: Articles count throughout the years based on the methodologies
Over the past decade, methods for detecting Self-Admitted Technical Debt (SATD) have changed significantly, as shown in Figure 4. There has been a clear shift from manual to more automated techniques, driven by a growing need for efficiency and accuracy in managing technical debt in software development.
From 2014 to 2016, SATD detection mainly relied on manual analyses. This process involved developers or analysts closely examining codebases and documentation to spot technical debt. While these manual methods provided valuable insights, they were time-consuming and susceptible to human error, highlighting the demand for more efficient solutions.
Between 2017 and 2019, the field began transitioning to semi-automated techniques. These methods combined manual efforts with automated tools to enhance the SATD detection process. The figure illustrates a gradual increase in the adoption of semi-automated approaches during this time, often utilizing regular expressions or Part-of-Speech (POS) tagging to reduce developer workload and improve consistency.
The period from 2020 to 2022 marked a significant move towards fully automated SATD detection methods. As depicted in the figure, there was a notable rise in the use of automated techniques, indicating the maturation of the field. These approaches employ advanced technologies such as natural language processing (NLP), machine learning (ML), and deep learning (DL) to automatically detect technical debt with minimal human input. As we move into 2023 and 2024, the trend toward automation in SATD detection continues to grow. The figure shows that automated methods have become the primary approach, outpacing manual and semi-automated techniques. This suggests that the future of SATD detection will likely focus on refining automated tools, possibly integrating more advanced AI and ML models.
As presented in Table 8, the methodologies are classified into three broad categories: Manual, Semi-Automated, and Automated approaches. These categories are determined based on the extent of human intervention and the utilization of NLP, ML, DL, Transformers models.
Table 8: Methodologies identified
Methodology Paper Reference NLP ML DL TRANSFORMERS
Manual No. 8, 37, 59, 39, 102 No. 8, 37, 59, 39, 102 - - -
Semi-Automated No. 36, 38, 40, 61, 41, 44, 45, 65, 46, 47, 49, 48 No. 36, 38, 40, 41, 44, 45, 46, 47, 49, 48 No. 65, 61 - -
Automated No. 50, 51, 52, 53, 54, 55, 56, 57, 77, 58, 79, 60, 62, 42, 43, 63, 64, 80, 81, 82, 83, 74, 84, 101, 85, 86, 87, 66, 88, 67, 89, 73, 68, 69, 76, 90, 91, 92, 70, 93, 94, 75, 95, 96, 99, 100, 71, 72, 97, 98 - No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 43, 64, 84, 101, 66, 67, 73, 68, 69, 70, 71, 72 No. 77, 78, 79, 62, 63, 80, 81, 82, 83, 74, 85, 86, 87, 88, 89, 76, 90, 91, 92, 94, 96 No. 93, 75, 95, 99, 100, 97, 98
Manual approaches significantly depend on human intervention, usually requiring the manual annotation of code comments or other textual software artifacts to detect SATD. Manual methods solely rely on NLP techniques. There is no application of ML, DL, or Transformers in the manual category, as these techniques are generally used for more automated processes. These methods are frequently time-consuming but offer high precision in identifying subtle or context-specific technical debt items. For example, Potdar and Shihab 8 conducted research using source code comments from several open-source projects. They manually classified comments as SATD or non-SATD based on specific comment patterns. Similarly, Maldonado and Shihab 37 employed manual labeling to construct a dataset for SATD that focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt. This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 51.
Semi-automated approaches combine manual review with automated tools to improve efficiency while retaining accuracy. Semi-automated methods predominantly rely on NLP, including part-of-speech (POS) tagging 36, and textual patterns 38; and a few of them use ML approaches 65, 61. However, there is no indication of the usage of Deep Learning or Transformers in these semi-automated techniques.
Automated approaches are characterized by the use of ML, DL, or Transformers and advanced NLP techniques to fully automate the SATD detection process. In order to detect SATD from issue tracking systems, Dai and Kruchten 50 created an automated method that makes use of an ML algorithm, namely of a Naïve Bayes (NB) classifier in addition to other NLP approaches like TF-IDF and TextRank. This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1-score. Recently, Li et al. 76 introduced a novel approach to automatically detect SATD from four different sources: source code comments, commit messages, pull requests, and issue tracking systems. The most recent trend in automated SATD detection includes the use of Transformers-based models (e.g., BERT 93, 75, RoBERTa 98, 95, and GPT 99).
Table 9: Textual artifacts Used. Studies that use more than one textual software artifact are highlighted with square brackets
Artifact Paper Reference NLP ML DL TRANSFORMERS
Code comment No. 8, 36, 37, 38, 51, 52, 53, 54, 55, 56, 57, 77, 58, 78, 40, 40, 60, 61, 62, 41, 42, 63, 64, 80, 81, 44, 82, 83, 45, 84, 101, 85, 86, 65, 66, 88, 67, 89, 73, 46, 68, 91, 92, 70, 93, 94, 75, 96, 99, 71, 49, 48, 72, 97,[69, 76, 90, 100] 8, 36, 37, 38, 40, 41, 44, 45, 46, 49, 48 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 64, 84, 101, 65, 66, 67, 73, 68,[69], 70, 71, 72 77, 78, 79, 62, 63, 80, 81, 82, 83, 86, 88, 89,[76], [90], 91, 92, 94, 96 93, 75, 99,[100], 97
Issue No. 50, 59, 39, 43, 102, 74, 87, 95, [69, 76, 90, 100] 59, 39, 102, 45 50, 84,[69] 74, 87,[76, 90] 95,[100]
Pull request No. 98,[69, 76, 100] - [69] [76] 98,[100]
Commit message No. 47,[69, 76, 100] 47 [69] [76] [100]
Moving forward from the methodological categorization, it is essential to understand the source or type of textual artifacts analyzed in the context of SATD. Table 9 categorizes the various studies on SATD detection based on the types of textual artifacts utilized. Code comments (such as from Java projects, blockchain projects, Dockerfile, and XML files), issues (from issue tracking systems and GitHub issue reports), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed. The application of NLP, ML, DL, and Transformers-based models across these artifacts shows how methodologies have evolved to handle different levels of complexity and scale.
Code comments are the most commonly analyzed artifact for SATD detection, with NLP, ML, DL, and Transformers models being widely employed (e.g., studies 8, 36, 93). While issues, pull requests, and commit messages are also explored, they are less frequently studied. Early research on SATD primarily relied on NLP to extract patterns from developer language in code comments and issues. However, more recent studies have incorporated ML, DL, and Transformers models (e.g., 76, 93, 98) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets. The increasing adoption of Transformers models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD. Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets. The increased use of pull requests and commit messages (e.g., 69, 76, 100) as sources of SATD detection in recent studies signals a growing recognition of these artifacts’ value in providing temporal and contextual data on technical debt. As the field moves forward, there is an increasing emphasis on automation and leveraging diverse textual artifacts, which aims to improve the scalability and reliability of SATD detection across varied software projects. The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
Answer to RQ2 Over the past decade, there has been a visible shift from manual towards fully automated approaches to improve efficiency and scalability in SATD detection. Additionally, the use of multiple textual artifacts has grown, enhancing the accuracy and robustness of SATD detection by integrating diverse sources from the software development process.
4.3 Results - RQ3: Effectiveness of the proposed approaches regarding their metrics evaluation results
To evaluate the effectiveness of the proposed SATD detection approaches, we examined the performance metrics reported in the studies included in our review. While a subset of the studies did not provide detailed performance results, the majority (49 out of 68) reported on key evaluation metrics such as precision, recall, and F1-score. The following section presents a detailed analysis of these metrics, utilizing boxplots to visualize the performance variability and trends across different SATD detection approaches.
Out of 68 relevant studies, 17 did not report their performance results, and 2 only provided AUC scores. Therefore, 49 out of the 68 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1-score. We collected all performance results presented in these 49 articles and organized them into a spreadsheet. Rather than focusing solely on the best-performing model from each study, we included the results from all model classifiers. This approach provides a more comprehensive and unbiased evaluation of SATD detection models 103. By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness. Limiting the analysis to only the best-performing models could lead to skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges. Moreover, analyzing all models enables a more robust cross-study comparison, allowing for the identification of consistent trends and patterns in model performance across different contexts, datasets, and techniques. This holistic approach strengthens the reliability of the conclusion and ensures their applicability across a wider range of scenarios, rather than being limited to cases of optimal performance.
We use boxplots to demonstrate the effectiveness of various approaches, as they are a useful graphical tool for showing variations between populations without relying on assumptions about data distribution 103. We compiled the effectiveness of various techniques as reported in the original papers, utilizing precision, recall, and F1-score derived from calculations based on the confusion matrix. Through boxplots, we showcase studies on precision, recall, and F1-score in relation to different factors of the models.
  • Precision = $\frac{TP}{TP+FP}$
  • Recall = $\frac{TP}{TP+FN}$
  • F1-score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
We acknowledge the potential for misinterpretation when aggregating the results of approaches across various studies without considering the diversity of datasets utilized. The rationale behind this methodological choice is to cultivate an inclusive understanding of each model’s performance and versatility across a wide range of contexts and scenarios. This inclusive approach enables the identification of universally effective strategies as well as those that excel under specific conditions, thereby providing a nuanced perspective on the potential applicability, hereby offering a holistic view of their effectiveness in detecting SATD. This methodology aligns with precedents in the field, notably reflected in the systematic literature review, such as the study of Hall et al. 103 and Malhotra 104. This precedent underscores the validity and relevance of cross-dataset comparisons in extracting meaningful insights that transcend the particularities of individual studies. By synthesizing findings across diverse datasets, this review aspires to contribute to the advancement of SATD detection approaches, offering valuable guidance for future research and practical applications in software engineering.
To avoid any misleading interpretations, we divide the process of SATD identification into two distinct tasks: SATD identification and SATD categorization. SATD identification focuses on determining whether a particular entity or instance is present in the text of a software development artifact, classifying it as either SATD or non-SATD. SATD categorization, on the other hand, takes this a step further by not only identifying the presence of SATD but also classifying it into specific types of technical debt. These categories may include labels such as code debt, design debt, requirement debt, and others, based on the linguistic elements found in the text.
In the following subsection, we present performance results for each SATD detection task, illustrated through boxplots.
4.3.1 Performance results of SATD identification
The box plot shown in Figure 5 provides a comparative overview of performance across different SATD identification methods, specifically for binary classification, which aims to accurately distinguish between SATD and non-SATD instances. The three key metrics displayed in the box plot—precision, recall, and F1-score—are standard measures for evaluating the effectiveness of classification models. Precision indicates the model’s ability to correctly identify instances as SATD or non-SATD, while recall measures how effectively the model captures all relevant SATD instances. The F1-score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other. Together, these metrics provide valuable insights into the performance of SATD detection methodologies in identifying technical debt within software projects.
Figure 5: Boxplot of SATD identification grouped by their approaches
The NLP-based approach demonstrates a relatively high precision of 0.79, indicating that it successfully identifies SATD in a significant number of cases when it classifies an instance as SATD. However, its recall is lower at 0.62, suggesting that it overlooks a considerable portion of SATD instances. The F1-score of 0.66 reflects the imbalance between precision and recall, highlighting that while NLP methods were foundational in early SATD detection, they now face challenges in fully capturing all relevant SATD instances.
In contrast, Machine Learning (ML) models offer a more balanced performance, achieving a precision of 0.73 and recall of 0.74, which results in an F1-score of 0.73. This improvement over NLP methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances. Historically, ML approaches have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
Deep Learning (DL) models further enhance performance, with precision and recall scores of 0.78 and 0.76, respectively, yielding a balanced F1-score of 0.76. This suggests that DL techniques are better at identifying SATD instances while minimizing false positives. The improvements in both precision and recall demonstrate the advancements deep learning architectures have brought to the complexities of SATD identification, enabling more accurate classification across diverse software contexts.
The Transformers-based approach achieves the best overall performance, with a precision of 0.83—the highest among all methods—and a recall of 0.76, leading to an F1-score of 0.76. These results underscore the superior capability of Transformers models to accurately identify SATD while retrieving a majority of relevant instances. As the latest advancement in SATD detection techniques, Transformers represent the current state-of-the-art, providing both the precision needed to avoid false positives and the recall required to capture a wide range of SATD instances.
Referring to Figure 5, we can further elaborate and illustrate the metric results in more detail, organizing them based on their respective models, as depicted in Figure 6. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
NLP-based models, particularly Textual Patterns and Annotation Tags, were assessed for their effectiveness. Textual Patterns (n=71) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives. However, its lower recall (0.62) and F1 score (0.66) suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification. In contrast, Annotation Tags (n=21) performed better in recall and F1 score, indicating a more balanced detection capability, although variability was present, as shown by the wider interquartile range.
ML models exhibited a broader range of performance. Among them, Random Forest (RF) (n=34) and Naive Bayes (NB) (n=37) provided the most consistent results, achieving high precision and F1 scores. This indicates their effectiveness in reducing false positives while capturing a majority of SATD instances. However, models like Support Vector Machine (SVM) (n=13) and k-NN (n=10) showed lower precision and recall, indicating challenges in accurately identifying SATD instances. Notably, Auto-sklearn and LightGBM performed well, though their small sample sizes (n=3 and n=1, respectively) limit the generalizability of their results.
DL models have made significant strides in SATD identification. BiLSTM (n=168), CNN (n=151), and LSTM (n=100) achieved high scores across all metrics, demonstrating consistent precision, recall, and F1 scores. These models effectively handle complex SATD detection tasks while minimizing false positives. However, models like Graph Neural Network (GNN) (n=35), Recurrent Neural Network (RNN) (n=20), and Graph Convolutional Network (GCN) (n=20) exhibited lower performance, particularly in recall, indicating they may miss important SATD instances. Models with smaller samples, such as Ladder Network (LN) (n=10) and Gated Graph Neural Network (GGNN) (n=10), also showed variability in performance, likely due to their limited test sizes.
Transformer-based models represent the most recent advancements in SATD identification. Among these, BERT (n=69), RoBERTa (n=39), and Transformer (n=40) achieved the highest precision, recall, and F1 scores. These models demonstrated remarkable consistency and reliability in identifying SATD instances, with minimal variability, making them leading techniques in SATD detection. Other Transformer models, such as ALBERT (n=18), DeBERTa (n=10), and GPT (n=3), also performed well, though slightly less robustly than the BERT-based models. DistilBERT (n=2) showed promise, but its limited sample size makes it difficult to draw definitive conclusions about its overall performance.
Figure 6: Models performance of SATD identification grouped by each approach
4.3.2 Performance result SATD categorization
Figure 7 offers insight into the precision, recall, and F1-score of these approaches, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled approaches differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others. The metrics provide insights into how effectively these approaches perform in a multiclass classification scenario, where distinguishing between different categories of SATD is more complex than a binary classification.
Figure 7: Boxplot of precision, recall, and F1-score of SATD categorization
NLP approaches demonstrate a low Recall of 0.11 and an F1-score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples. This limitation affects their effectiveness in real-world applications, particularly in complex multiclass classification tasks. In contrast, ML approaches show notable improvement, achieving a Precision of 0.75 and a Recall of 0.57, resulting in a balanced F1-score of 0.52. These results suggest that ML models provide better accuracy in SATD classification, although they still miss a considerable number of true positive cases, as indicated by the Recall.
DL approaches perform well across all metrics, with a strong Precision of 0.77 and a high Recall of 0.76. However, the F1-score of 0.62 suggests that while they successfully capture most true positives, there is still potential for improvement in balancing false positives and false negatives. Interestingly, Transformer-based approaches, despite being more advanced, exhibit slightly lower overall performance in this instance, with a Precision of 0.69, Recall of 0.59, and an F1-score of 0.54. While transformers typically excel in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on their performance in this specific SATD identification task. Nonetheless, transformers still outperform NLP approaches and are comparable to ML models in this context.
Referring to Figure 7, we can provide a comprehensive evaluation of specific types of SATD categorization, organized by their respective models, as shown in Figure 8. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
Figure 8: Models performance of SATD categorization grouped by each approach
In the NLP group, two approaches, Textual patterns (n=46) and Annotation tags (n=2), were evaluated. The Textual Patterns approach showed highly variable performance, with precision hovering around 0.6, while the recall and F1 scores dropped significantly, indicating that while this method can accurately categorize SATDs in some cases, it struggles to consistently retrieve all SATD instances. The large interquartile range and the presence of outliers highlight the inconsistency of this method in different contexts. In contrast, Annotation tags, although evaluated on only two samples, showed relatively stable and high precision, recall, and F1 scores, indicating strong performance for the multiple cases tested, albeit with limited generalizability. These findings confirm the limitations of NLP techniques in handling the complexities of SATD categorization. While effective in some cases, particularly for high precision, NLP methods often miss critical SATD instances, leading to suboptimal recall and overall lower F1-scores. This aligns with the review’s conclusion that NLP approaches, although foundational in early SATD detection, are no longer sufficient for comprehensive and consistent categorization of SATD.
ML models showed a wide range of results, reflecting the diversity in algorithm complexity and adaptability to SATD categorization. Random Forest (RF) (n=114) and Auto-sklearn (n=108) showed the highest performance, consistently providing high precision and recall, indicating that these models are able to balance false positives and missed SATD examples effectively. However, Naive Bayes (n=17) showed weaker performance, particularly in terms of recall, indicating that it may miss important SATD cases. Other models, such as SVM (n=25), XGBoost (n=50), and Logistic Regression (n=24), showed lower recall and more significant performance variation across the dataset, reflecting their challenges in categorizing SATD consistently.
DL models showed significant improvement over the NLP and ML approaches, specifically in their ability to maintain high precision, recall, and F1 scores across larger data sets. CNN (n=163), CNN-BiLSTM (n=20), and LSTM (n=70) stood out for their consistent and robust performance, with higher recall rates indicating their effectiveness in detecting a wider range of SATD events without significantly increasing false positives. GGNN and Feedforward neural network (FFNN), despite their smaller sample sizes, also showed strong performance, although the small dataset makes it difficult to draw broader conclusion about their generalizability.
Transformers-based models, specifically BERT (n=47), ALBERT (n=24), and RoBERTa (n=13), were the highest performing approaches in SATD categorization. BERT emerged as the best performer, with high precision, recall, and F1 scores across all evaluations, demonstrating its ability to accurately and consistently detect SATDs with minimal false positives and missed events. ALBERT and RoBERTa also performed strongly, although evaluated on smaller sample sizes, their high precision and F1 scores reinforced their value in SATD detection.
From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1-score. It is important to note that Figure 6 represents SATD identification aimed at distinguishing SATD from non-SATD (binary classification), while Figure 8 illustrates the results of SATD categorization of specific SATD types (multiclass classification). To select the best-performing algorithm from the boxplot, we primarily focused on the F1-score (dark blue bars) as it offers a balance between precision and recall. We identified the algorithm with the highest median F1-score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance suggesting more consistent performance. Additionally, we examined the precision and recall (light and mid-blue bars) to ensure a balance between the two, prioritizing algorithms with high values in both. In addition, we minimized the influence of algorithms with numerous outliers, as they indicate inconsistent performance, even if their median scores were high. Finally, we factored in the n value (number of experiments), giving preference to algorithms tested more frequently, as a higher n suggests more reliable and stable results across different trials.
Table 10: Summary of highest performance results based on SATD detection tasks
Approach Metric Evaluations
Precision Recall F1-score
SATD identification
NLP Annotation tags Annotation tags Annotation tags
ML SMO LR SMO
DL GNN BiLSTM BiLSTM
Transformers GPT Transformer BERT
SATD categorization
NLP Annotation tags Annotation tags Annotation tags
ML SMO NB SMO
DL GGNN GGNN GGNN
Transformers BERT BERT BERT
The best-performing algorithms for each approach based on the different SATD types is presented in Table 10. In both the SATD identification and categorization tasks, BERT consistently emerges as the best-performing model. For SATD identification, BERT’s high median F1-score, tested across a large number of experiments (n=69), showcases its strong reliability and generalizability. Its low variance and minimal outliers further demonstrate its consistent and stable performance across different datasets. Leveraging its advanced Transformer-based architecture, BERT excels at capturing the context and nuances in SATD instances, making it highly effective in balancing precision and recall.
Similarly, in the SATD categorization task, BERT outperformed other models, achieving the high median F1-score with a significant number of experiments (n=47). Its minimal variance and absence of outliers reinforce its reliability and consistency in this task as well. Altogether, these factors position BERT as the most robust and effective model for both SATD identification and categorization, making it the top choice in each task.
Answer to RQ3 BERT (a Transformers-based model) emerges as the top performer in both SATD and non-SATD identification and the categorization of specific SATD types, providing balanced precision, recall, and F1 scores. SMO (an ML model) and GGNN (a DL model) also show strong performance within their respective domains, with SMO excelling in precision and F1 score, while GGNN leads in identifying specific SATD types. Overall, Transformer models, particularly BERT, represent the most advanced and effective approach to SATD detection.
4.4 Results - RQ4: Proposed tools for SATD detection
To address this research question, we conducted a thorough review of existing tools developed over the past decade. While the detection and management of SATD have been extensively studied, there remains a noticeable gap in translating research into practical tools for developers. This section examines tools specifically designed for SATD detection, offering insights into advancements in automation and machine learning, as well as the diverse approaches used to address various forms of technical debt.
Table 11 presents a collection of tools proposed for SATD identification across different software projects. This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, machine learning integration, and varied strategies for managing different types of technical debt. Below, we describe these tools and their significance in the broader context of SATD identification, emphasizing key trends that have emerged over the last decade.
Table 11: SATD detection tools proposed
Name Reference Year Task Type Description Status
eXcomment No. 36, 48 2015 Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
SATD Detector No. 55 2018 Identification A Java library and an Eclipse plug-in. The tool automatically identifies whether comments are SATD or non-SATD and integrates with an IDE to help manage these comments Obsolete
DebtHunter No. 64 2021 Identification, Categorization A machine learning-based approach for detecting self-admitted technical debt Available
FixMe No. 72 2021 Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
SATDBailiff No. 67 2022 Identification A tool designed to mine, identify, and track SATD Available
A browser extension No. 74 2022 Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of a rOpenSci R package Not available
DebtViz No. 90 2023 Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
The tools proposed for SATD identification reflect a diverse range of approaches to addressing technical debt in software projects. eXcomment 36, 48, and SATD Detector 55 focus on identifying both SATD and non-SATD comments, with eXcomment integrating into an IDE for real-time detection. A browser extension 74 leverages machine learning to detect various SATD types, including documentation and usability debt, highlighting the potential of ML for broader SATD detection. SATDBailiff 67 provides comprehensive tracking and monitoring of SATD over time, while FixMe 72 is a GitHub bot specifically designed to alert developers about unresolved, “on-hold SATD" in repositories. Lastly, DebtViz 90 offers a more complete solution, combining SATD detection, classification, visualization, and monitoring in one platform, emphasizing the need for visual tools that make technical debt more understandable to stakeholders. Together, these tools demonstrate an increasing focus on automation, monitoring, and comprehensive SATD management, though the overall number of tools remains limited relative to the research interest in the field.
This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 68 relevant studies, only 7 articles proposed specific tools for SATD identification. This disparity points to a concerning gap between the significant research interest in SATD and the actual development of practical, automated solutions that can assist developers in managing technical debt. While a range of tools has been introduced, the overall lack of available tools indicates that much of the research has yet to translate into actionable, industry-ready solutions. Addressing this gap will be critical for advancing the practical management of technical debt in software projects, moving from theoretical frameworks toward robust, scalable automation.
Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, machine learning integration, and addressing different types of technical debt. However, despite significant research interest, only a few tools have been developed, revealing a gap between theoretical advancements and the availability of practical, automated solutions for managing technical debt in real-world software projects.
4.5 Results - RQ5: Primary challenges associated with SATD detection
To address the challenges associated with SATD detection, we conducted a thorough analysis of the relevant literature, focusing on the difficulties commonly highlighted by researchers. These challenges reflect the complexities inherent in identifying SATD across various software projects and methodologies.
We identified nine key challenges related to SATD identification extracted from the relevant studies, as summarized in Table 12, with relevant studies emphasizing each challenge. These challenges were determined by examining sections commonly located in the ‘Discussion’ or ‘Threat to Validity’ sections of the papers reviewed. We extracted all discernible challenges, whether implicitly or explicitly mentioned, associated with SATD detection. Each challenge is explained in this section and further elaborated on below.
Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets
No Challenge Paper Reference
1 Error-proneness dues to manual analysis No. 8, 38,[37, 36, 50, 53, 59, 39, 49]
2 Imbalanced data No. [54, 77, 78, 40, 79, 60, 61, 62, 64, 81, 82, 83, 45, 85, 86, 88, 68, 93, 94, 96, 71], 56, 63, 80, 66, 89, 46, 47, 92, 75, 95, 97
3 Generalizability No. [50, 53, 59, 39, 37, 55, 77, 40, 60, 61, 64, 78, 81, 45, 85, 86, 88, 68, 93, 94, 96, 49, 99, 71], 58, 74, 101, 87, 65, 69
4 Explainability No. [77, 78, 83, 85, 86, 88, 99, 44, 100]
5 False negatives No. [54, 79, 62]
6 Risk of reproducibility No. [ 44, 100, 67]
7 Computational costs No. [44, 82, 83, 67, 100]
8 Overlap of features among debt types No. [37, 66]
9 Tagging and comment relevance No. [36, 49]
Manual Analysis
The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 38. Since the process relies on human judgment, it becomes prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD differently 36. This subjectivity in analysis can lead to varying outcomes in tasks like pattern recognition and feature engineering, making the process less reliable.
The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can severely impact the effectiveness of machine learning models that rely on this data for training. When models are trained on datasets that contain inaccuracies or inconsistencies, their predictive accuracy and generalization capabilities suffer. This challenge becomes even more pronounced when scaling up to larger datasets, where consistent and precise labeling is critical to maintaining the quality of the training data. Without robust and reliable analysis, the models may struggle to handle the increased complexity of larger data, leading to reduced performance and misidentification of SATD instances.
Imbalanced data
In most software projects, SATD comments represent only a small portion of the total code comments, resulting in a significant class imbalance. This imbalance can adversely affect the performance of SATD identification models, as the models may tend to focus on non-SATD comments, which dominate the dataset. As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 66.
This imbalance frequently leads to a high number of false negatives 105, where the model fails to detect SATD instances, undermining its effectiveness. Since machine learning models generally find it difficult to learn patterns from underrepresented classes 106, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics. Consequently, the model’s ability to identify and classify SATD instances diminishes, particularly when faced with a large volume of non-SATD items. Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that the models are not biased toward detecting the majority class while neglecting the crucial SATD instances.
Generalizability
One of the primary challenges associated with the identification of SATD is the issue of generalizability. Generalizability refers to the ability of SATD identification models to maintain their accuracy and effectiveness when applied across different software projects and domains. This challenge arises due to the diverse nature of software projects, which often exhibit significant variations in coding practices, domain-specific terminologies, and architectural designs 44.
Models trained on data from one specific project tend to be tailored to the characteristics of that project, limiting their performance when applied to other projects with different practices. For example, the way SATD is expressed in comments may differ across projects due to differences in the development team’s language, domain jargon, or code style conventions 58, 39. These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when faced with unfamiliar coding practices or new domains.
Explainability
Another primary challenge in the SATD identification process, particularly in machine learning and deep learning-based models, is the lack of explainability. This issue arises because many advanced models, such as neural networks and deep learning architectures, function as “black boxes," meaning their internal decision-making processes are not easily interpretable by users 107. These models generate predictions without providing transparent explanations about the underlying reasoning for classifying an instance as SATD, non-SATD, or even specific types of SATD.
The absence of explainability presents a significant obstacle to the adoption and trustworthiness of SATD detection tools. Moreover, explainability is crucial for validating the accuracy of SATD identification models. If developers cannot trace the logic behind a model’s predictions, it becomes difficult to diagnose errors or refine the model to reduce false positives or false negatives.
False negatives
False negatives are another concern in SATD detection. They occur when SATD items are incorrectly classified as non-SATD, leading to undetected technical debt in the codebase. This happens when models fail to identify instances where developers have either implicitly or explicitly admitted technical debt 38. As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 12, 108.
Risk of reproducibility
Another primary challenge in the context of machine learning and supervised models-based models is the issue of reproducibility 44, 67. These models often involve complex architectures with numerous parameters that must be fine-tuned for optimal performance. However, many studies fail to provide sufficient details on the hyperparameters, training configurations, or the specific processes used to achieve their results. Additionally, the lack of publicly available code exacerbates the reproducibility problem, as other researchers are unable to replicate the exact experimental setup. This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these approaches.
Computational costs
Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly deep learning approaches. These models require significant resources during training, including high-performance GPUs, large memory capacities, and extended computation time. This high computational cost creates several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
Additionally, the extended computation time required for training deep learning models presents a challenge in terms of practicality. SATD identification models often need to be retrained or fine-tuned to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices. In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of deep learning models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
Overlap of features among debt types
Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 66. This overlap makes distinguishing between different debt types difficult, leading to potential misclassification. The ambiguity in distinguishing between SATD and other debt types can reduce the accuracy of identification models. If features that are common to multiple debt types are not carefully handled, models may struggle to correctly identify SATD.
Tagging and comment relevance
Finally, tagging and comment relevance is a critical challenge in identifying SATD. Developers often use tags such as “TODO, “FIXME," or “XXX" in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes. While these task-related tags are commonly employed as indicators of SATD, they are not always relevant to technical debt. In many cases, tags like “TODO" or “FIXME" can represent a simple reminder or a non-technical debt issue, thereby creating noise in the identification process. Moreover, not all SATD instances are explicitly marked by such tags, further complicating the identification process.
Answer to RQ5 We identified nine major challenges in detecting SATD, including issues like manual analysis errors, imbalanced data, lack of generalizability, and explainability in advanced models. Other challenges include false negatives, reproducibility risks, computational costs, overlapping features among debt types, and the relevance of tagging and comments, all of which hinder effective SATD identification.
5 Discussion
The findings from this systematic review reveal several key trends and challenges in the detection of SATD over the past decade, demonstrating both the evolution and fragmentation within the field. SATD detection approaches have progressed significantly, shifting from simple keyword-based NLP methods to more complex ML, DL, and Transformers-based models. While each advancement has brought increased performance and automation, the development of practical and scalable SATD detection tools remains limited.
5.1 Evolution of SATD detection techniques
Early SATD detection techniques primarily relied on basic NLP methods, such as keyword matching and textual patterns, to identify technical debt in code comments. These approaches, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts. From 2017 onward, ML techniques such as Naive Bayes (NB), Support Vector Machines (SVM), and Random Forests became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives. Despite this progress, these methods still faced challenges related to generalizability across different datasets and projects, as well as difficulty in explaining the rationale behind the classification of certain comments as SATD.
The rise of deep learning methods from 2020 to 2022 marked a significant leap forward in SATD detection. Models like Long Short-Term Memory (LSTM) networks and Convolutional Neural Networks (CNN) enabled the identification of more complex patterns in developer comments and project artifacts. These approaches improved recall and reduced false negatives, offering a more balanced approach to identifying SATD across diverse datasets. However, the high computational cost of training and deploying these models remains a major barrier to widespread adoption.
Most recently, Transformers-based models, such as BERT and RoBERTa, have emerged as the leading approaches for SATD detection. These models have demonstrated superior performance in capturing the context and semantics of developer comments, enabling more accurate and generalizable SATD identification. The ability of Transformers to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection. However, issues related to the explainability of these models and their scalability for industrial use still persist.
5.2 Challenges in SATD detection
Despite the advancements in SATD detection techniques, several challenges hinder the practical adoption of these methods in real-world software development environments. A key challenge in SATD detection is the imbalance in datasets—SATD instances are often much less frequent than non-SATD instances, which negatively impacts model performance. This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, are far less represented, further complicating detection. Although several techniques have been employed to address data imbalance in SATD detection, future research should explore more advanced methods, such as leveraging large language models (LLMs) to better handle imbalanced data and improve model performance.
Generalizability remains another key challenge. Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their ability to perform well on new or proprietary codebases. This lack of generalizability reduces the practical utility of these models in diverse software development environments, where coding practices, project documentation, and domain-specific language can vary significantly. The explainability of advanced models, particularly deep learning and Transformer-based approaches, is also a concern. Without a clear understanding of why certain comments are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows. Future work should focus on developing explainable AI techniques for SATD detection to ensure that these models’ decision-making processes are transparent and understandable to developers. Research on enhancing the interpretability of DL and Transformers models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
Another significant barrier to adoption is the high computational cost associated with training and maintaining deep learning and Transformers-based models. While these models offer improved accuracy, their resource-intensive nature makes them difficult to implement for smaller development teams or in environments with limited computational resources. This challenge highlights the need for more efficient algorithms or alternative approaches that can deliver high accuracy without the associated computational burden.
5.3 Lack of tools for SATD detection
Despite the significant academic interest in SATD detection, the translation of research into practical tools remains limited. This review identified only seven tools developed specifically for SATD detection over the past decade, reflecting a notable gap between theoretical research and practical application. While tools like SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, the overall number of available tools is insufficient to meet the needs of diverse software development teams. Additionally, the tools that do exist often lack the ability to integrate seamlessly into modern development environments, limiting their utility for real-time SATD management.
The limited development of practical tools also highlights the need for further collaboration between academia and industry. Researchers must focus on creating tools that not only demonstrate high performance in academic settings but also address the specific needs and constraints of industrial software development, such as scalability, usability, and integration with existing development environments. Bridging the gap between theoretical advancements and practical, scalable tools is essential for achieving widespread adoption in industry. Future research should prioritize developing user-friendly tools that can be integrated into continuous integration (CI) environments, enabling real-time SATD detection and providing actionable insights for developers. Moreover, these tools should facilitate collaboration within teams working on large-scale software projects, ensuring that they are both practical and effective in real-world development scenarios.
5.4 Future Directions
To address the challenges identified in this review, future research should prioritize enhancing the generalizability, explainability, and efficiency of SATD detection models. Techniques such as transfer learning, which allow models to be trained on one dataset and then applied to others with minimal retraining, could significantly improve generalizability. At the same time, incorporating explainable AI (XAI) techniques is crucial to ensure that the decision-making processes of these models are transparent and understandable for developers. Developing more interpretable models, or employing post-hoc explanation techniques, could foster trust in these tools and encourage their adoption in real-world projects. Further research on enhancing the interpretability of deep learning and Transformer models—potentially through visualizations or simplified intermediate representations—would empower practitioners to better understand, trust, and effectively utilize these models in software development settings.
There is also a need for the development of more practical SATD detection tools that can be easily integrated into existing development environments. These tools should focus on providing real-time feedback to developers, enabling them to address technical debt as part of their normal workflow. By bridging the gap between research and practice, the software engineering community can ensure that SATD detection techniques contribute meaningfully to the long-term maintenance and evolution of software systems.
6 Threats to Validity
As with any systematic literature review (SLR), the results of this study are subject to various validity threats. These are categorized as internal, external, and construct validity threats, along with mitigation strategies to minimize their impact.
6.1 Internal validity
Internal validity refers to the reliability of the methodology employed in the study. One key threat to internal validity in this SLR is related to the formulation of the search strings. In the pilot search, several iterations of search strings were tested, starting with a more general search (e.g., “self-admitted technical debt" OR SATD and NLP), which led to a broader, less focused set of results. Over time, more refined strings were developed, including terms like “detection" and “identification," as well as the use of boolean operators and the wildcard (*) for broader term variations.
The final search string (S3), which combined terms such as “self-admitted technical debt" OR SATD with NLP OR “natural language processing" AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies. This iterative process helped mitigate the risk of missing key studies during the initial literature search.
Furthermore, data extraction inaccuracies could threaten the reliability of our findings. To reduce this risk, the data extraction process was double-checked, and disagreements were resolved through discussions between the authors. A pilot data extraction was also conducted to further minimize potential bias.
6.2 External validity
External validity refers to the extent to which the findings of this SLR can be generalized to other contexts beyond the studies reviewed. A primary threat to external validity in this research is the limited generalizability of the results due to the scope of the reviewed datasets. All of the studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints faced in proprietary or large-scale industrial environments.
Open-source datasets often have characteristics, such as well-documented codebases or community-driven development practices, that differ significantly from proprietary software projects, which may involve stricter deadlines, unique domain requirements, or higher security and compliance constraints. As a result, some of the trends and findings related to SATD detection may not extend to these more complex and regulated environments. For example, the performance of machine learning or deep learning models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
6.3 Construct validity
Construct validity pertains to whether the study accurately measures the concepts it aims to investigate. A significant threat in this SLR is the comparison of results from different studies, which can be misleading if the studies use different methodologies or report metrics inconsistently. Directly comparing the precision, recall, and F1-scores of models trained on heterogeneous datasets without considering their context could lead to inaccurate conclusion.
To address this, we employed a grouping strategy to compare similar studies and ensured that the datasets were organized before any comparisons were made. By contextualizing the results within groups, we aimed to make the comparisons more meaningful. Furthermore, this methodology is based on established practices within the field, notably exemplified by Hall et al. 103 and Malhotra 104, which similarly undertakes the task of comparing results across different studies to deduce broad-based insights into algorithmic performance in varying contexts.
7 Conclusion
This SLR provides an in-depth analysis of SATD detection approaches over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated machine learning and deep learning models. Among the key findings, it is evident that while NLP-based techniques were foundational in early SATD detection, they struggle with recall and overall performance in comparison to ML and DL approaches. Transformer-based models, particularly BERT, emerged as the most effective in capturing complex patterns in textual artifacts, offering superior performance across multiple metrics.
However, several challenges persist, including issues related to dataset diversity, the scalability of models, and the explainability of AI techniques. The reliance on open-source datasets in many studies limits the generalizability of the findings, and more research is needed to adapt these models to varied industrial environments. Moreover, the imbalance of SATD datasets continues to affect the robustness of model performance, and integrating multiple textual artifacts such as commit messages and pull requests presents further complexity. Furthermore, the lack of widely adopted tools for SATD detection in industry highlights a gap between research advancements and practical applications.
In conclusion, while the field of SATD detection has advanced considerably, there remains a need for further research and collaboration between academia and industry to ensure these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
[…]
*Data availability Supplementary material related to this article, including all the boxplots as images, can be found online at https://github.com/edisutoyo/satd-detection-slr
[…]
*Declaration of competing interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Financing of Higher Education (BPPT) and Indonesia Endowment Fund for Education (LPDP).

### v4

@@
 Abstract
-[Abstract] Technical debt (TD) represents the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals.
+[Abstract] Technical debt (TD) refers to the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals.
 Self-Admitted Technical Debt (SATD) occurs when developers explicitly acknowledge these trade-offs in the codebase, typically through comments or annotations.
-Automated detection of SATD has become an increasingly important research area, particularly with the rise of natural language processing (NLP), machine learning (ML), and deep learning (DL) techniques that aim to streamline SATD detection.
-This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and 2024, focusing on the evolution of techniques from NLP-based models to more advanced ML, DL, and Transformers-based models such as BERT.
-The review identifies key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1-score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and the explainability of models.
-The findings suggest that while early NLP methods laid the foundation for SATD detection, more recent advancements in DL and Transformers models have significantly improved detection accuracy.
-However, challenges remain in scaling these models for broader industrial use.
-This SLR offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
+SATD detection has become an increasingly important research area, particularly with the rise of learning-based techniques that aim to streamline SATD detection.
+This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and early 2025, focusing on the evolution of techniques from heuristic-based techniques to more advanced ML, DL, and Transformer-based models.
+It examines key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1 score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and explainability.
+The findings reveal that while early heuristic-based techniques laid the foundation for SATD detection, more recent advancements in DL and Transformer models have significantly improved detection accuracy.
+However, challenges remain in scaling these models for broader industrial adoption.
+This review offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
 […]
@@
 1 Introduction
-Business needs often require developers to be able to deliver new products and features as rapidly as possible, in order to reach the market.
-In fact, a developer must be able to balance carefully designing applications and seamlessly coding their features.
-In the business or corporate world, however, schedules and limited resources frequently hinder developers from generating perfect code before deploying it into products 1.
-This becomes commonplace when stakeholders cut several technical processes to catch up on momentum, especially in startup practices 2, 3.
-A tight deadline frequently leads to technical debt, as developers lack time to consider long-term solutions 4.
-Technical debt (TD) is often used to describe the trade-off that occurs when an organization (or a developer) chooses to prioritize the delivery of new features or functionality over the long-term maintenance of the software 5.
-TD can occur in any software development project and can be caused by a variety of factors, including lack of resources, time, knowledge, or experience on the part of the development team 6, 7.
+Business needs often require developers to rapidly deliver new products and features to reach the market quickly.
+It is important for developers to carefully balance design and implementation to seamlessly add features.
+However, in the business or corporate world, schedules and limited resources often hinder developers from writing perfect code before incorporating it into products 1.
+This becomes commonplace when stakeholders cut technical processes to gain momentum, especially in startups 2, 3.
+Tight deadlines often lead to technical debt because developers lack time to consider long-term solutions 4.
+Technical debt (TD) is often used to describe the trade-off that occurs when an organization prioritizes delivering new features or functionality over maintaining the software in the long term 5.
+TD can occur in any software development project and may be caused by a variety of factors, including insufficient resources, time, knowledge, or experience on the part of the development team 6, 7.
 Among various types of TD, self-admitted technical debt (SATD) has gained considerable attention due to its explicit acknowledgment within the source code comments by developers.
-SATD represents instances where developers intentionally or unintentionally leave behind suboptimal design or implementation choices, often with a clear understanding of the resultant negative impact on the codebase 8.
+SATD refers to instances in which developers, either intentionally or unintentionally, make suboptimal design or implementation choices, despite having a clear understanding of the negative impact these choices will have on the codebase 8.
 This acknowledgment is not merely an admission of suboptimal practices but a vital piece of information that can guide future development activities.
 Recognizing SATD is crucial for maintaining software quality and sustainability, as it can accumulate over time and impede future development efforts.
-Over the past decade, many approaches to detecting SATD have emerged, driven by advances in natural language processing (NLP), machine learning (ML), and deep learning (DL).
-These approaches aim to automate the detection of SATD, facilitate their management, and helping software teams address it proactively.
-Alongside these approaches, various methodologies have been proposed to detect SATD in different stages of the software development life cycle, leveraging manual and semi-automated, and automated techniques.
+Over the past decade, various approaches to SATD detection have emerged, evolving from heuristic-based techniques to more advanced machine learning (ML), deep learning (DL), and Transformer-based models.
+These approaches aim to automate SATD detection, facilitate its management, and enable software teams to address it proactively.
+Alongside these approaches, various methodologies have been proposed to detect SATD, ranging from manual to semi-automated to automated methodologies.
 However, the diversity of techniques, methodologies, tools, and evaluation metrics used in SATD detection has led to a fragmented landscape, making it challenging for researchers and practitioners to gain a holistic understanding of the field.
@@
 Time frame and scope.
-Our review is the most extensive to date, covering relevant studies published between 2014 and 2024.
+Our review spans a decade-long period from January 2014 to April 2025, analyzes 74 studies on SATD detection, and provides the most up-to-date synthesis of SATD research.
 In contrast, the review by Sierra et al.
-focuses on a narrower time frame, covering studies published between 2014 and 2018 9.
+9 focuses on a shorter interval (2014–2018) and captures earlier SATD detection efforts.
+Similarly, Gu et al.
+10 conduct a lightweight review that covers literature until 2022, but primarily focuses on evaluating detection models across different software artifacts rather than providing a broad literature synthesis.
 Systematic approach.
-We followed the original and rigorous procedures outlined by Kitchenham and Charters 10 for conducting systematic reviews.
-In contrast, Sierra et al.’s study 9 appears to be based on a survey approach, without explicitly detailing how studies were sourced.
+We follow the established and rigorous procedures outlined by Kitchenham and Charters 11 for conducting systematic reviews.
+In contrast, Sierra et al.
+9 appears to be based on a survey approach, without explicitly detailing how studies are sourced.
+Gu et al.
+10 conduct an empirical study, but their review methodology does not follow an SLR framework, as it primarily assesses the generalization of SATD models across programming languages and software artifacts.
 Comprehensiveness.
-In addition to relying on search engines, our review included a bidirectional snowballing process to ensure we captured all relevant articles. As a result, we analyzed a broader range of studies.
-Profoundness.
-Unlike Sierra et al.
-9, who did not report the metric evaluation results, our review includes detailed analysis of metrics such as precision, recall, and F1-score, categorizing them based on the approaches and methodologies used.
-Analysis.
+In addition to relying on digital library searches, our review employs a bidirectional snowballing process to capture all relevant studies and results, resulting in 74 analyzed papers and making our review broader and more representative of the field.
 While Sierra et al.
-9 broadly discussed SATD activities without focusing deeply on detection approaches, our review provides a more detailed evaluation of detection techniques, their effectiveness, and the tools proposed for SATD detection.
-In this study, we conduct and evaluate a systematic literature review that focuses on the detection of SATD, particularly using NLP-based approaches and their combinations.
-The objectives and key contributions of our study are as follows:
-  • We consolidate and categorize various approaches employed to detect SATD, revealing various techniques and identifying critical trends over the past decade.
-  • We examine methodologies proposed for SATD detection, showcasing their impact on automation and accuracy.
-  • We investigate tools developed for SATD detection over the past decade, providing an overview of their effectiveness.
-  • We compare the best-performing approaches by evaluating precision, recall, F1-score, and generalizability across different software systems and datasets.
-  • We identify primary challenges in SATD detection, offering insights for future research and improvement.
-In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section  3 describes the methodology utilized to conduct the SLR.
+9 also employ snowballing, their review covers a narrower time frame (2014–2018), which may limit the discovery of additional relevant studies.
+In contrast, Gu et al.
+10 do not explicitly perform snowballing, primarily focusing on evaluating existing SATD detection models rather than conducting a comprehensive literature search.
+Depth of analysis.
+Prior reviews, such as Sierra et al.
+9 and Gu et al.
+10, do not consistently synthesize reported findings.
+Our review extends this by examining both methodological characteristics (e.g., degree of automation) and empirical evidence (e.g., reported performance metrics), offering a more integrated perspective on how SATD detection techniques evolve and perform.
+In this study, we conduct a systematic literature review on SATD detection, evaluating existing approaches and their effectiveness. The objectives and key contributions of our study are as follows:
+  • We consolidate and classify SATD detection approaches, tracing the evolution from heuristic-based methods to ML, DL, and Transformer-based models.
+This taxonomy provides researchers with a structured view of methodological progress in the field.
+  • We compare methodologies in terms of their level of automation (manual, semi-automated, and automated) and synthesize reported performance metrics (precision, recall, and F1 score) across diverse datasets and software systems.
+This synthesis provides a consolidated overview of the techniques and highlights the best-performing models reported in prior studies.
+  • We conduct the first systematic review of SATD detection tools, assessing their availability, capabilities, and reported effectiveness.
+This analysis reveals persistent gaps between research prototypes and practical tool adoption in industry settings.
+  • We identify and synthesize nine recurring challenges in SATD detection (e.g., dataset imbalance, model generalizability and explainability, and adoption barriers), which highlight key opportunities for future research.
+In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section 3 describes the methodology utilized to conduct the SLR.
 Sections 4 and 5 summarize and discuss the obtained results, while Section 6 identifies the threats to validity of this study.
@@
 2 Background
-This section introduces the concepts underpinning this research, related to NLP and self-admitted technical debt, as well as the relevant works.
+This section introduces the key concepts underlying this study, focusing on self-admitted technical debt (SATD) and the role of natural language processing (NLP) techniques that are directly relevant for analyzing the textual artifacts used in SATD detection.
 2.1 Self-Admitted Technical Debt (SATD)
-The term “technical debt” was coined by software developer Ward Cunningham.
+The term “technical debt” (TD) was coined by software developer Ward Cunningham.
 He initially used the metaphor for communicating to non-technical stakeholders the importance of budgeting resources for refactoring 5.
@@
 In other words, there is a trade-off between delivering high-quality, effective, and optimized code and the short-term benefits of achieving release dates 6, 4.
-A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 11, 12.
-The term “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 13.
+A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 12, 13.
+The “debt” metaphor further implies that, just as financial debt accrues interest, the longer suboptimal code remains in the system, the more costly and time-consuming it becomes to remove.
+This emphasizes the compounding nature of technical debt over time, where delayed repayment leads to greater maintenance burdens and reduced productivity.
+The concept “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 14.
 It represents the debt incurred when development teams opt for quick and easy solutions in the short term, risking long-term negative impacts.
-The term “TD” encompasses below-par aspects of the software that developers are aware of but lack resources to address, such as outdated documentation, unperformed tests, complex code needing refactoring, and known defects left unattended 14, 15.
-These immature artifacts can lead to challenges in meeting quality standards and cause unforeseen delays in implementing necessary changes.
-Numerous empirical studies have recently concentrated on examining source code comments and other textual artifacts that indicate technical debt, often known as self-admitted technical debt (SATD) 8.
-The source used to detect SATD is usually in the form of natural language often expressed through text, such as code comments, commit messages, or other textual artifacts; therefore, SATD detection significantly relies on NLP-based approaches.
-2.2 Natural Language Processing (NLP)
-Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language, encompassing tasks like speech recognition, natural language understanding, and natural language generation 16, 17.
-Chowdhury 17 explains that the objective of NLP researchers is to understand how people perceive and use language so that the appropriate tools and techniques can be developed to assist computers in understanding and using natural languages to complete the necessary tasks.
-NLP deals with processing natural language, which is the language used by humans in text or audio form.
-Computational linguistics, rule-based human language modeling, statistical models, machine learning, and deep learning are combined to enable computer programs to interpret and process human language 18, 19.
-In the context of SATD detection, plain text is used as input, including project artifacts such as source code comments, commit messages, pull requests, and issue tracking systems 20, 21.
-Source code comments serve as annotations within the code, providing documentation directly incorporated by programmers 22.
-These comments are essential for understanding a system and may contain important information, such as the purpose of code additions, progress status on collective tasks, or reasons for specific changes 23.
-Comments can express concerns that need attention in the future (e.g., the TODO keyword), identify emerging issues, and guide future actions 24, 25.
-They provide valuable context specific to their authors and are human-readable.
-Recently, automated approaches utilizing NLP have gained traction in addressing software engineering problems, including the detection of TD, which is called SATD detection.
-NLP has been effectively used to automatically identify the presence of TD and even categorize specific types of SATD.
-Notably, numerous studies have integrated NLP with machine learning (ML) models or deep learning (DL) architectures and also Transformers, aiming to enhance model performance (evidence is provided in Subsection 4.1).
+TD encompasses various suboptimal aspects of software that developers acknowledge but lack the immediate resources to address.
+These may include outdated documentation, missing tests, complex code in need of refactoring, and known defects left unresolved 15, 16.
+Such deficiencies can degrade software quality over time, making maintenance more challenging and causing unforeseen delays when implementing necessary changes.
+In recent years, numerous empirical studies have focused on identifying technical debt through textual artifacts, particularly source code comments and other developer-written text, a practice commonly referred to as self-admitted technical debt (SATD) 8.
+The primary sources for SATD detection include code comments, commit messages, issue tracker discussions, and other natural language-based software artifacts.
+As SATD is expressed in natural language, its detection relies heavily on language processing techniques to analyze and classify technical debt instances.
+Therefore, SATD detection can be regarded as an application of NLP within the broader context of software engineering, as it relies heavily on natural language techniques to identify and classify SATD.
+2.2 NLP in SATD Detection
+Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language.
+It includes tasks such as speech recognition, natural language understanding, and natural language generation 17, 18.
+Within the scope of SATD detection, NLP is applied in a more focused way as it provides the means to transform developer-authored textual artifacts such as code comments, commit messages, issue tracker discussions, and pull requests into structured features that automated models can analyze 19, 20.
+For instance, code comments often contain explicit markers of pending work (e.g., TODO) or implicit expressions of design compromise, which can be mined using NLP-based approaches 21, 22, 23.
+Unlike traditional static code analysis approaches that detect proxies of technical debt, such as code smells or rule violations, SATD detection aims to uncover technical debt as explicitly acknowledged by developers themselves 8, 24.
+This makes SATD a unique and valuable complement to other TD detection methods, as it directly reflects developers’ awareness of design trade-offs and postponed tasks 25.
+In prior studies, SATD detection has been approached from two complementary perspectives.
+Some research focuses on determining whether a given software artifact contains any indication of debt 24, 26, thereby providing an overview of its extent within a system.
+Other research takes this further by distinguishing between different forms of debt, such as design, architecture, requirement, or test debt, based on the linguistic patterns found in the text 24, 27, 28.
+This finer-grained view is particularly useful because different types of SATD have distinct implications for maintenance and repayment strategies 9.
+Together, these perspectives highlight how detection approaches can support both the recognition and the deeper understanding of technical debt in practice 29.
+Over time, SATD detection techniques have evolved in line with advances in NLP, moving from simple keyword-based heuristics to classical machine learning classifiers trained on TF-IDF features, and more recently to deep learning and Transformer-based models that leverage semantic embeddings 24, 30, 31.
+Although these approaches demonstrate significant progress, the field is still maturing and continues to face open questions that require systematic investigation.
+This underlines the importance of consolidating and critically examining the state of the art, which is the focus of this review.
+While the background section introduced SATD and its detection techniques, a systematic investigation is required to consolidate existing research findings.
+To achieve this, we employ a structured literature review methodology.
+The following section outlines our research design, including our research questions, study selection process, and data extraction methods, ensuring a rigorous and reproducible review.
 3 Research Methodology
-This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al26, 27, 10.
-The procedures for SLR are broken down into three stages: planning, conducting, and reporting the SLR 10.
-We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure  1).
-Figure 1: Graphical representation of the SLR process used in this article
+This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al.
+32, 33, 11.
+The procedures for SLR are divided into three stages: planning, conducting, and reporting 11.
+We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure 1).
+Figure 1: Graphical representation of the SLR process used in this article.
 The review process comprises several steps.
 Initially, research questions were formulated for the SLR.
-Subsequently, a search strategy was devised, involving the identification of search strings and the selection of online databases for locating primary studies.
+Subsequently, a search strategy was devised, involving the development of search strings and the selection of online databases to locate primary studies.
 The third stage encompassed the identification of pertinent studies aligned with the research questions, along with the establishment of inclusion and exclusion criteria.
-Following this, bidirectional snowballing was conducted utilizing references from the source articles.
-After that, we created a data extraction form to answer the research questions.
+To ensure comprehensive coverage, a bidirectional snowballing approach was employed.
+This involved examining the references cited in the source articles and exploring articles that cited these sources, using Google Scholar [^1] https://scholar.google.com as the primary tool, as recommended in prior guidelines 34.
+Subsequently, we developed a data extraction form to address the research questions.
 A data extraction form was then crafted to address the research questions.
 Lastly, a data synthesis approach was developed to analyze and summarize the research findings.
-Figure 2 depicts the article selection procedure for determining the primary studies used in this SLR. The following subsections provide detailed information about the selection process.
-Figure 2: An overview of the steps in determining primary studies
 3.1 Research Questions
-Our SLR includes all relevant research outlining the various SATD detection techniques.
-To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 28.
+Our SLR reviews all relevant research outlining the various SATD detection techniques.
+To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 35.
 Boolean operators were applied to structure the search effectively: OR was used to include synonyms and alternative terms, while AND linked the key terms, as shown in Table 1.
 This approach ensured a comprehensive and focused search strategy.
-Table 1: PICOC Criteria
+Table 1: PICOC Criteria.
 Parameter Description Keywords Used
@@
 Intervention Approach or method for dealing with a certain problem “nlp” AND “technical debt”
-Comparison Not aplicable Not aplicable
+Comparison Not applicable Not applicable
 Outcome Related to factors of importance to practitioners “detection” OR “identification” OR “prediction”
 Context Context in which intervention take places “software development” OR “software engineering”
-According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect self-admitted technical debt.
-Therefore, the aim of this SLR is to answer the following questions:
-RQ1: What approaches have been employed to detect self-admitted technical debt (SATD) in software development?
-Rationale: This question aims to understand the progress made in detecting SATD. By surveying existing studies, we can detect trends and potential future research directions.
-RQ2: What methodologies have been proposed to detect SATD in software development?
-Rationale: This question aims to analyze the various approaches and tools developed over the past decade for SATD detection.
-RQ3: How effective are these approaches regarding their metrics evaluation results?
-Rationale: This question aims to evaluate the performance and practicality of the identified approaches by comparing their precision, recall, and F1-score.
+According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect SATD.
+Therefore, this SLR aims to answer the following questions:
+RQ1: What types of techniques have been employed to detect SATD in software development, and how do these techniques handle different types of textual artifacts?
+Rationale: This question aims to categorize and analyze the various types of techniques that have been proposed for SATD detection across different types of textual artifacts, providing insight into trends and gaps in the field.
+RQ2: What levels of automation have been proposed in methodologies for detecting SATD in software development, and to what extent do these methods rely on human interventions across different types of textual artifacts?
+Rationale: This question focuses on understanding the extent to which SATD detection methods incorporate automation, highlighting the progress and identifying gaps in automating the detection process across different types of textual artifacts.
+RQ3: Which SATD detection techniques are reported as the most effective based on evaluation metrics?
+Rationale: This question aims to review and synthesize the reported precision, recall, and F1 scores of the identified techniques to provide an overview of their observed performance.
+The goal is not to perform an independent benchmark but to consolidate how techniques have been reported in the literature and to identify which are most frequently reported as achieving the highest effectiveness.
 RQ4: What tools have been proposed for SATD detection?
@@
 3.2 Search Protocol
-Before beginning the SLR planning stage, it is necessary to conduct a pilot search on a database such as Scopus to observe if there are any SLRs with the same research theme.
-There is no need to conduct a new SLR if there are existing SLRs on the same topic 29.
-In this systematic review of the literature, no results covering the topic of the capabilities of the NLP approach in identifying technical debt were discovered; therefore, the SLR can be conducted as new research.
+Before initiating the SLR planning stage, it is essential to conduct a pilot search in databases such as Scopus [^2] https://www.scopus.com to determine whether existing SLRs have already addressed the same research theme.
+If a relevant SLR is found, conducting a new one may not be necessary 36.
+Our preliminary search revealed no existing SLRs encompassing the breadth of approaches for detecting SATD.
+Consequently, this SLR fills a critical gap in the literature and offers a timely contribution to the field.
 The SLR’s search strategy consists of three stages: pilot search, primary search, and snowballing.
-In the pilot search, multiple search strings are used on online databases to determine the most effective one.
-The primary search uses the identified search string to find relevant literature in online source databases.
-The search concludes with a snowball search 30.
-The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore, ScienceDirect, ACM Digital Library, and Springer.
-These databases, as proved by previous software engineering SLRs 31, 32, 33, 34, provide an excellent overview of the software engineering literature and are trustworthy enough to cover our scope of the literature.
+In the pilot search, multiple search strings are used in online databases to identify the most effective.
+The primary search uses the identified search string to retrieve relevant literature from online databases.
+The search concludes with a snowball search 37.
+The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore Digital Library, ScienceDirect, ACM Digital Library, and Springer.
+These databases, as demonstrated by previous software engineering SLRs 38, 39, 40, 41, provide an excellent overview of the software engineering literature and are sufficiently comprehensive to cover our scope.
 Moreover, we employed the “Advanced Search” feature in these databases to obtain full-text and metadata results.
-Table 2: General search strings utilized during pilot search
+The first author communicated the results of a pilot search using the search string S1 to the second author.
+The second author randomly selected 10% of the articles to check their relevance.
+Based on the results, the second author proposed modifying the search string to S2 because the results from S1 remained too general.
+Similar to the initial process, the iteration was repeated until it was decided to use the final search string, namely S3 (see Table 2).
+Table 2: General search strings utilized during pilot search.
 String ID Search String
-S1 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”)
-S2 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
-S3 (“self-admitted technical debt" OR SATD) OR (“technical debt" AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering" OR “software development")
-The results of a pilot search using search string S1 were communicated to the second author by the first author.
-The second author selected 5% articles at random to see if they were relevant.
-The second author proposed altering the search string to S2 because the search string S1 result was still too general.
-Similar to the first process, iteration is repeated until it is decided to utilize the final search string, namely search string S3.
-As illustrated in Table 2, through a pilot search with multiple search strings, the most effective one (i.e., S3) was selected.
-The search string was designed with synonyms, alternative keywords, and boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies.
-Drawing from the pilot search conducted earlier, for this search string S3, we incorporated three supplementary terms commonly associated with tasks related to SATD identification: detect, identify, and predict.
-Additionally, wildcard asterisk (*) was used to broaden the search for variations of words 35.
-The final search string (i.e., S3) was applied to the mentioned 6 online databases to find relevant primary studies for the research questions as described in Subsection 3.1.
-We utilized the “Advanced Search” feature in these databases to acquire full-text and metadata results.
-Table 3: Search results from selected databases
+S1 (“self-admitted technical debt” OR SATD) AND (NLP OR “natural language processing”)
+S2 (“self-admitted technical debt” OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
+S3 (“self-admitted technical debt” OR SATD) OR (“technical debt” AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering” OR “software development”)
+Through a series of pilot searches with multiple search strings, the most effective one (i.e., S3) was selected.
+The search string was designed with synonyms, alternative keywords, and the Boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies.
+Building on the earlier pilot search, we expanded search string S3 by adding three supplementary terms that are commonly used in SATD identification tasks: detect, identify, and predict.
+Additionally, the wildcard asterisk (*) was used to broaden the search for variations of words 42.
+The final search string (i.e., S3) was then applied to the six selected online databases to identify relevant primary studies addressing the research questions described in Subsection 3.1.
+We utilized the “Advanced Search” feature in these databases to obtain full-text and metadata results, as summarized in Table 3.
+Table 3: Search results from selected databases.
 No Database Search Result
-1 Web of Science 51
-2 Scopus 1,004
-3 IEEE Xplore 168
-4 ACM Digital Library 41
-5 Springer 125
-6 ScienceDirect 51
-Total 1,440
-As depicted in Figure 2, the initial step involved searching multiple online databases using predefined search queries to collect a comprehensive set of studies relevant to the research topic.
-As shown in Table 3), a total of 1,440 studies were retrieved at this stage.
-To manage such a large dataset, both authors worked collaboratively to remove duplicates, which resulted in 1,119 unique records.
-Following the duplicate removal, a filtering process based on metadata by using inclusion and exclusion criteria presented in Table 4 was conducted.
-At this stage, the titles, abstracts, and keywords of the remaining studies were examined to determine their relevance to SATD detection.
-A full-text analysis of these 69 studies was then performed to evaluate the quality and relevance of the research in greater detail.
-The first author initially reviewed the metadata and conducted an independent assessment of the full papers, focusing on their contributions to SATD detection approaches.
-After completing the initial evaluation, the first author discussed the findings with the second author, who served as a supervisor throughout the process.
-This discussion provided a critical review of the selection criteria and the individual assessments.
-To ensure that no bias influenced the selection process, the second author independently reviewed a subset of the full-text studies.
-These results were then compared with the first author’s evaluations, and any discrepancies or differences in judgment were carefully discussed and resolved.
-This collaborative approach minimized the risk of subjective bias and ensured that each study was rigorously and objectively evaluated based on its quality and relevance.
-To further ensure consistency in the evaluation process, the test-retest strategy outlined by Kitchenham and Charters 10 was employed, rather than relying on inter-rater agreement.
-This strategy involved discussions between the first and second authors to maintain a consistent evaluation framework.
-Additionally, the second author cross-checked the data extraction results, which were shared in a spreadsheet [^1] The spreadsheet is available in the shared dataset area at https://github.com/edisutoyo/satd-detection-slr.
-Any disagreements that arose during the process were addressed through collaborative discussions until consensus was reached.
-After this comprehensive review and resolution of all discrepancies, both authors agreed to select 63 relevant studies from 6 online databases to be included in the systematic review.
-This thorough, unbiased, and transparent process ensured the inclusion of only the most relevant and high-quality studies.
-3.3 Article Selection and Snowballing Process
-Inclusion and exclusion criteria must be precisely defined for study selection.
-The inclusion criteria should include an automated approach to detect technical debt.
-Consequently, the primary studies were chosen based on comparing titles, keywords, abstracts, and full texts (i.e., introduction, methodology, results, and conclusion).
-We employed 6 inclusion criteria and 5 exclusion criteria, as shown in Table 4):
-Table 4: Inclusion and exclusion criteria
+1 Web of Science 56
+2 Scopus 1,118
+3 IEEE Xplore 184
+4 ACM Digital Library 49
+5 Springer 181
+6 ScienceDirect 69
+Total 1,659
+3.3 Study Selection
+As shown in Figure 2, we selected primary studies through a staged screening procedure to retain only relevant and sufficiently detailed evidence for this review.
+The initial search across multiple digital libraries returned 1,659 records.
+After duplicate removal, 1,245 unique records remained.
+We then applied a three-stage screening process guided by the inclusion and exclusion criteria in Table 4.
+Figure 2: An overview of the steps in determining primary studies.
+First, we performed a metadata screening of titles, abstracts, and keywords to assess topical relevance, which reduced the set to 194 studies.
+Second, we screened studies for research methodology and eligibility (e.g., whether the study reported an SATD identification approach and provided sufficient methodological detail), yielding 85 studies.
+Third, we conducted full-text screening to confirm relevance and assess the study’s contribution and reporting quality, yielding 69 studies focused on SATD detection.
+These 69 studies constitute the initial core set of primary studies for our SLR.
+Table 4: Inclusion and exclusion criteria.
 No Inclusion criteria
@@
 3 For research that has both publication types (journal and proceeding version), only the journal version will be selected
-4 The study must discuss approaches for identifying and measuring the performance of SATD detection
-5 The study must be published between the years 2014 and 2024
+4 The study must discuss techniques or approaches for identifying and measuring the performance of SATD detection
+5 The study must have been published between 2014 and April 2025
 6 The study reported must be related to SATD in software engineering or software development
@@
 5 All other secondary and tertiary studies
-To further enhance the comprehensiveness of the review, a bidirectional snowballing technique 30 was employed.
-Both forward and backward snowballing were performed, meaning that the references cited in the selected studies, as well as the studies that cited them, were examined to identify any additional relevant research that might have been overlooked in the initial search.
-This process resulted in the inclusion of 5 additional studies.
-The second author actively contributed to the snowballing process, assessing the relevance and quality of newly identified studies.
-In the final stage, the 63 studies selected from the full-text filtering process were combined with the 5 studies identified through bidirectional snowballing, resulting in a total of 68 primary studies.
-This collaborative effort between the authors ensured that the selection process was both thorough and objective, ultimately leading to a comprehensive review of the existing literature on SATD detection approaches.
+To ensure the reliability of our inclusion and exclusion process, we followed the recommendations of Kitchenham and Charters 11, who suggest that consistency of study selection can be verified through a test-retest procedure as an alternative to inter-rater agreement.
+The first author conducted the initial screening of studies across all stages (metadata, research methodology, and full-text analysis).
+To check the consistency of the labeling process, the first author also applied a test-retest approach by re-evaluating a random sample of the studies.
+In addition, to minimize potential bias, a subset of 15% of the studies at each filtering stage was independently reviewed by the second author.
+Both validation steps yielded full agreement with the initial decisions, and no disagreements were found, indicating that the risk of misclassification was negligible.
+The replication package, including the screening spreadsheet and supplementary material, is publicly available.
+After completing these validation steps, both authors agreed to select 69 relevant studies from six online databases for inclusion in the systematic review.
+This rigorous and transparent process ensured the inclusion of only the most relevant and high-quality studies.
+The replication package, including the screening spreadsheet and supplementary material, is publicly available. [^3] https://github.com/edisutoyo/satd-detection-slr
+3.4 Snowballing Process
+To improve the completeness of the review, we employed bidirectional snowball sampling in accordance with Wohlin’s guidelines 37.
+In line with the practical recommendations of Brereton et al.
+34, we used Google Scholar to support both backward and forward snowballing.
+Backward snowballing was performed by manually screening the reference lists of the included studies, whereas forward snowballing was conducted by identifying and screening papers that cited those included studies.
+For backward snowballing, references within the files of the selected studies were carefully examined.
+Each reference was assessed based on its title and abstract to determine its relevance.
+When necessary, the full texts of the referenced studies were retrieved and reviewed for further evaluation.
+Initially, the collected references were filtered by their titles and abstracts to assess alignment with the research questions.
+Subsequently, the full texts of potentially relevant studies were analyzed to determine whether they met the inclusion criteria.
+To ensure the reliability of the snowballing process, both authors were involved in assessing the relevance of the identified studies.
+The first author initially screened the references and citations, and the second author independently reviewed the candidate papers and cross-checked their eligibility.
+Any disagreements were discussed until a consensus was reached.
+Only one candidate paper required clarification at this stage, which was resolved through discussion.
+This indicates that disagreements were minimal and resolvable.
+This rigorous process helped uncover additional studies that might have been missed during the initial search.
+Five additional studies were included as a result.
+In the final stage, the 69 studies selected through full-text screening were combined with the five studies identified through bidirectional snowballing, yielding a total of 74 primary studies.
+This collaborative effort ensured that the selection process was both thorough and objective, thereby leading to a comprehensive review of the existing literature on SATD detection approaches.
 3.5 Quality Assessment Criteria
-After finalizing the list of articles for the SLR, we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions.
-The quality of the selected articles was assessed using the questions presented in Table 5, following the procedure recommended by Kitchenham 26.
-All 68 relevant articles we discovered passed the quality evaluation, and thus, our SLR relies on these 68 primary studies.
-Table 5: Quality assessment criteria
+After finalizing the list of articles for the systematic literature review (SLR), we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions.
+We assessed the quality of the selected articles using the questions in Table 5, following the procedure recommended by Kitchenham 32.
+All 74 articles deemed relevant satisfied the predefined quality assessment criteria.
+Accordingly, our SLR is based on these 74 primary studies.
+Table 5: Quality assessment criteria.
 No Questions
@@
 3 Does the paper describe a research methodology?
-4 Does the article employ an algorithm or model for SATD detection?
+4 Does the article employ an algorithm or technique for SATD detection?
 5 Does the study provide research findings and answer the research questions?
-6 Does the article report metric evaluation performance or at least report results of the approach?
+6 Does the article report metric evaluation performance, or at least report results of the approach?
 7 Does the article clearly state the dataset used?
-All of the listed articles were published between January 2014 and June 2024.
-Our searches for articles were completed by the end of June 2024, supplemented by articles obtained through bidirectional snowballing.
+All of the listed articles were published between January 2014 and April 2025.
+Our searches for articles were completed by early May 2025 and were supplemented by articles identified through bidirectional snowballing.
 As a result, it is highly unlikely that we missed any published articles during our period due to publication delays.
@@
 We determined the required information and attributes for data summarization to address our research questions.
-This facilitated the data extraction process, where we read the full text of each article and extracted the relevant information based on specified attributes.
+This facilitated data extraction, in which we read the full text of each article and extracted the relevant information based on specified attributes.
 To answer the research questions mentioned in Subsection 3.1, we extracted data items listed in Table 6 from each selected study.
 The extracted data was then entered into a spreadsheet for further analysis.
-Table 6: Data extracted from each study
+As part of the extraction, we also examined the evaluation setup adopted by each study.
+Since our review specifically focuses on within-project evaluations (e.g., k-fold cross-validation or random splits performed on the same dataset or project), we excluded cross-project or out-of-distribution evaluations from the performance synthesis in RQ3.
+This decision was made to ensure comparability of results across studies and to avoid conflating performance metrics obtained under substantially different evaluation settings.
+Table 6: Data extracted from each study.
 No Data Name Description Relevant RQ
@@
 6 Algorithms used Algorithm or model used in the study RQ1, RQ2, RQ4
-7 TD Types The types of TD, e.g., design or requirement TD R2, RQ3
-8 Number of Class Number of class(es)/TD types used R2, RQ3
-9 Dataset used Data source and metadata of dataset used R1, RQ4
-10 Metrics Evaluation used Types of metric evaluations used for performance measurement RQ4
-11 Performance Evaluations Performance measurement results using metric evaluations RQ4
+7 Tool proposed Name or description of any tool proposed in the study RQ1, RQ2, RQ4
+8 TD types The types of TD, e.g., design or requirement debt RQ2, RQ3
+9 Number of class Number of class(es)/TD types used RQ2, RQ3
+10 Textual artifacts used Data source or software artifacts used RQ1, RQ4
+11 Metrics evaluation used Types of metric evaluations used for performance measurement RQ3
+12 Performance summary All included studies report within-project validation (e.g., k-fold CV, random splits). Cross-project results are not synthesized in this review RQ3
 13 Challenges Issues or challenges related to SATD detection RQ5
+To ensure the reliability of the data extraction and categorization process, we followed the guidance of Kitchenham and Charters 11.
+As a ‘gold-standard procedure’, they recommend that data extraction should be performed independently by at least two reviewers whenever feasible.
+Because full dual independent extraction for all included studies was not feasible under our resource constraints, we applied two complementary reliability checks: (i) a test–retest procedure in which the first author re-extracted and re-categorized a random sample after an interval to assess decision stability, and (ii) an independent cross-check by the second author on 15% of the extracted items.
+[^4] This conforms to the guidelines as suggested in section 6.4.3 of 11: ‘Alternatively, a test-retest process can be used where the researcher performs a second extraction from a random selection of primary studies to check data extraction consistency’.
+The shared spreadsheet used to take these decisions is available in the supplementary material at https://github.com/edisutoyo/satd-detection-slr) These checks are in line with the guidelines of Kitchenham and Charters 11, and resulted in full agreement with the initial decisions, and no disagreements were observed.
+For transparency and replicability, the complete dataset and categorization results are made available in our replication package.
+With the methodology in place, we present our findings from the 74 selected studies.
+The results are organized according to our research questions, beginning with an analysis of the techniques used for SATD detection, followed by an evaluation of automation levels, effectiveness metrics, available tools, and key challenges in this domain.
 4 Results
-This section addresses the posed research questions. To that purpose, the outcomes of the 68 primary studies are presented and discussed.
-4.1 Results - RQ1: Approaches used for detecting SATD in software development
-As part of the results for RQ1, we categorized the relevant studies into three main groups based on their approach: papers using NLP, those employing machine learning (ML), and studies utilizing deep learning (DL) or Transformers (see Figure 3 and Table 7).
-Any disagreements that arose during the grouping process were resolved through discussion.
-Figure 3 shows the number of papers published by approach: NLP, ML, DL, and Transformers.
-Over the past decade, the identification of SATD has evolved significantly, driven by advancements in NLP and ML techniques.
-Early methods relied primarily on traditional NLP techniques, focusing on keyword detection and textual patterns in code comments.
-As depicted in Figure 3, NLP dominated initially, with limited exploration of advanced computational methods.
-From 2017 to 2019, ML approaches were increasingly introduced to improve SATD detection accuracy.
-Between 2020 and 2022, the use of DL and Transformer architectures surged, marking a shift toward more sophisticated, data-driven techniques.
-These models captured complex patterns in software artifacts, improving SATD detection.
-As of 2023 and 2024, DL and Transformer models continue to lead SATD detection efforts.
-Figure 3: Articles count throughout the years based on their approaches
-Table 7: Approach identified. Studies that fall into more than one approach are highlighted with square brackets
-Approach Paper Reference Technique examples
-NLP No. 8, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49 POS tagging, Regular Expressions, Textual patterns, etc.
-Machine Learning No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64 65, 66, 67, 68, 69, 70, 71, 72,[73, 74, 75, 76] NB, ME, SVM, kNN, etc.
-Deep Learning No. 77, 78, 79, 80, 81, 82, 82, 83, 84, 85, 86, 87, 88, 89 90, 91, 92, 93, 94,[73, 74, 75, 76] CNN, LSTM, BiLSTM, etc.
-Transformers No. 95, 96, 97, 98, 99, 100,[73, 74, 75] BERT, ALBERT, GPT, etc.
-NLP approaches have played a fundamental role in detecting SATD.
-These methods generally concentrate on text preprocessing tasks that encompass stemming, lemmatization, and stopwords removal.
-The objective is to process and evaluate textual data from code comments or documentation, where technical debt is frequently acknowledged by the authors themselves.
-As presented in Table 7, other methods include manual analysis and Parts-of-Speech (POS) tagging 36, 40, 41, 49, which aid in classifying and identifying relevant information.
-Regular expressions 38, 46, keyword-labeled 61, and textual patterns 45 are commonly employed to semi-automate the identification of certain technical debt indicators in large codebases.
-As the complexity of software systems increased, so has the demand for more sophisticated approaches in technical debt identification, resulting in the use of ML classifiers.
-These algorithms include Naïve Bayes (NB) 50, 55, 59, Maximum Entropy (ME) 51, 67, Support Vector Machines (SVM) 52, and k-Nearest Neighbor (kNN).
-Logistic Regression (LR) 101 and Random Forest (RF) 56, 43 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt.
-Furthermore, XGBoost 66, 68, a relatively new addition to the ML classifier, improves performance by refining gradient boosting approaches.
+This section addresses our research questions and discusses the findings from the 74 primary studies included in this review.
+4.1 Results - RQ1: Techniques employed for detecting SATD in software development across textual artifacts
+To classify the relevant studies, we followed a bottom-up categorization process.
+During data extraction, we recorded the algorithms and models used in each primary study (see Table 6) and grouped them into higher-level categories based on commonalities observed in the literature and prior surveys.
+This process led to four main categories (see Table 7 and Figure 3): (i) heuristic-based approaches (rule-based, keyword, or patterns matching methods), (ii) machine learning approaches (classical supervised models such as SVM, NB, and RF), (iii) deep learning approaches (neural architectures such as CNN, LSTM, BiLSTM, and GNN), and (iv) Transformer-based approaches.
+While Transformers are technically a specialized form of deep learning, we treat them as a distinct category because they represent a clear methodological shift: unlike recurrent models such as LSTM, Transformers rely on self-attention mechanisms 43, are more scalable, and have become the dominant paradigm in recent SATD studies.
+This separation follows trends observed in both NLP and SE literature 44, 45, where Transformers are increasingly treated as a distinct family of models due to their state-of-the-art performance and widespread adoption.
+Moreover, recent large language models (LLMs) are built on the Transformer architecture 46, further reinforcing the distinctiveness of this category.
+Table 7: Technique identified. Studies that fall into more than one technique are highlighted with square brackets.
+Technique Paper Reference Technique examples
+Heuristic No. 8, 47, 25, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60 Annotation tags, Textual patterns matching, etc.
+Machine learning No. 26, 24, 61, 62, 63, 64, 65, 66, 67, 27, 68, 69, 70, 71, 72 73, 28, 74, 75, 76, 77, 78, 79, 80, 81,[82, 83, 84, 30] NB, ME, SVM, kNN, etc.
+Deep learning No. 85, 86, 87, 88, 89, 90, 90, 91, 92, 93, 94, 95, 96, 97, 81 98, 99, 100, 101, 102,[82, 83, 84, 30, 31, 103, 81] CNN, LSTM, BiLSTM, etc.
+Transformer No. 104, 105, 106, 107, 108, 10, 103, 109,[82, 83, 84, 31] BERT, ALBERT, ChatGPT, etc.
+The categorization process followed the reliability procedures described in Subsection 3.6, which included test-retest and cross-checking by the second author.
+Only one disagreement was recorded and resolved through consensus, indicating that the grouping was stable and consistent.
+Figure 3 presents the distribution of published papers categorized by their detection techniques, including heuristic-based methods, ML, DL, and Transformer-based approaches.
+Over the past decade, various approaches to SATD detection have emerged, evolving from heuristic-based methods to more advanced ML and DL techniques.
+Early methods relied primarily on traditional heuristic techniques, focusing on keyword detection and textual patterns in code comments.
+As depicted in Figure 3, heuristic methods initially dominated, with limited exploration of advanced computational methods.
+From 2017 to 2019, ML techniques were increasingly introduced to improve SATD detection accuracy.
+Between 2020 and 2022, DL and Transformer architectures surged, marking a paradigm shift toward data-driven techniques capable of capturing more complex semantic patterns.
+From 2023 onward, DL and Transformer models became the predominant approaches, with Transformers slightly surpassing DL in 2024 and maintaining their momentum into early 2025.
+This trend highlights a growing reliance on Transformer-based models for advancing SATD detection.
+Figure 3: Distribution of articles by technique over the years.
+Heuristic-based techniques have long played a fundamental role in detecting SATD.
+These approaches typically rely on lightweight, rule-driven text processing and pattern matching rather than model training.
+As summarized in Table 7, heuristic pipelines may also incorporate manual inspection and Parts-of-Speech (POS) tagging 47, 50, 51, 59 to support the categorization of debt-related expressions.
+In addition, regular expressions 48, 56, keyword-labeled patterns 69, and other textual pattern rules 55 are widely used to semi-automate the detection of recurring SATD indicators in large codebases.
+As software systems have grown in complexity, the need for more advanced technical debt detection methods has intensified, driving the adoption of ML classifiers to enhance accuracy and scalability.
+These algorithms include Naïve Bayes (NB) 26, 64, 27, Maximum Entropy (ME) 24, 74, Support Vector Machines (SVM) 61, and k-Nearest Neighbor (kNN).
+Logistic Regression (LR) 110 and Random Forest (RF) 65, 53 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt.
+Furthermore, XGBoost 28, 75 and LightGBM 75, relatively new additions to the ML classifier, improve performance by refining gradient boosting approaches.
 The rise of DL has introduced even more advanced techniques for detecting SATD.
-Convolutional Neural Network (CNN) 77, 83, 85, 90, Long Short-Term Memory (LSTM) 79, 81, and Bidirectional Long Short-Term Memory (BiLSTM) 78, 82 are among the most prominent techniques used.
+Convolutional Neural Network (CNN) 85, 91, 93, 98, Long Short-Term Memory (LSTM) 87, 89, and Bidirectional Long Short-Term Memory (BiLSTM) 86, 90 are among the most prominent techniques used.
 They are particularly useful in processing structured data and identifying patterns that are not easily captured by traditional ML algorithms.
-Deep learning approaches have been extensively studied and are shown to be more critical in automatically detecting technical debt.
-Transformers are the most recent development in the detection of SATD.
-By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 95, 96, 100, 97, A Lite BERT (ALBERT) 98, Robustly Optimized BERT (RoBERTa) 98, and DistilBERT 98, have transformed the field of NLP.
-Transformers excel at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt.
-The literature reflects a growing trend towards using these models to achieve higher accuracy and more robust identification processes, especially in dealing with large and complex datasets typical of modern software systems.
-Answer to RQ1 Four main approaches have been identified to detecting SATD: traditional NLP, ML, DL, and Transformers.
-Over time, the focus has shifted from basic NLP techniques to more advanced methods like ML and DL, with transformers now leading the way, due to their effectiveness in capturing complex textual patterns.
+DL techniques have been extensively studied and have been shown to be more effective for automatically detecting technical debt.
+The transformer is the most recent development in SATD detection.
+By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 104, 105, 10, 106, 31, A Lite BERT (ALBERT) 107, Robustly optimized BERT approach (RoBERTa) 107, DistilBERT 107, DistilRoBERTa 109, as well as more recent large-scale models such as Flan-T5 103 and ChatGPT 108, have transformed the field of NLP.
+Transformer excels at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt.
+The literature reflects a growing trend toward using these models to achieve higher accuracy and more robust identification processes, particularly for large, complex datasets typical of modern software systems.
+Table 8: Textual artifacts used. Studies that use more than one textual software artifact are highlighted with square brackets.
+Artifact Paper Reference Heuristic ML DL Transformer
+Code comment No.
+8, 47, 25, 48, 24, 61, 62, 63, 64, 65, 66, 85, 67, 86, 50, 50, 68, 69, 70, 51, 52, 71, 72, 88, 89, 54, 90, 91, 55, 92, 110, 93, 94, 73, 28, 96, 74, 97, 82, 56, 75, 99, 100, 77, 101, 102, 84, 105, 108, 78, 59, 80, 103, 58, 79, 106, 81,[76, 30, 98, 10, 31] 8, 47, 25, 48, 50, 51, 54, 55, 56, 59, 58 24, 61, 62, 63, 64, 65, 66, 67, 68, 69, 52, 72, 92, 110, 73, 28, 74, 82, 75,[76], 77, 78, 79, 80 85, 86, 87, 70, 71, 88, 89, 90, 91, 94, 96, 97,[30], [98, 31], 99, 100, 102, 105, 103, 81 101, 84, 108,[10, 31], 106, 103, 81
+Issue No. 26, 27, 49, 53, 111, 83, 95, 104, 109, [76, 30, 98, 10, 31] 27, 49, 111, 55 26, 92,[76] 83, 95,[30, 98, 31] 104, 109,[10, 31]
+Pull request No. 107,[76, 30, 10, 31] - [76] [30, 31] 107,[10, 31]
+Commit message No. 57, 60,[76, 30, 10, 31] 57, 60 [76] [30, 31] [10, 31]
+Building on the approach categorization, it is essential to understand the sources or types of textual artifacts analyzed in the context of SATD.
+Table 8 categorizes the various studies on SATD detection based on the types of textual artifacts utilized.
+Code comments (such as from Java projects 24, 85, blockchain projects 59, 77, Dockerfile 55, 10, XML files 92, 10, and quantum software projects 60), issues (from issue tracking systems 30, 31 and GitHub issue reports 104, 109), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed.
+The application of heuristic, ML, DL, and Transformer-based models across these artifacts shows how approaches have evolved to handle different levels of complexity and scale.
+Code comments are the most commonly analyzed artifact for SATD detection, with heuristic, ML, DL, and Transformer-based models being widely employed (e.g., studies 8, 47, 101).
+While issues, pull requests, and commit messages are also explored, they are less frequently studied.
+Early research on SATD primarily relied on heuristic-based techniques to extract patterns from developer language in code comments and issues.
+However, more recent studies have incorporated ML, DL, and Transformer models (e.g., 30, 101, 107, 109, 108) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets.
+The increasing adoption of Transformer models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
+Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD.
+Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets.
+The increased use of pull requests and commit messages (e.g., 76, 30, 10, 31) as sources for SATD detection in recent studies signals a growing recognition of their value in providing temporal and contextual data on technical debt.
+As the field advances, there is an increasing emphasis on automation and on leveraging diverse textual artifacts to improve the scalability and reliability of SATD detection across varied software projects.
+The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
+Our synthesis reveals a clear research bias, with most SATD detection studies focusing heavily on code comments, whereas other artifacts, such as issue trackers, commit messages, and pull requests, remain underexplored.
+This concentration on comments is understandable given their accessibility, yet it limits the diversity and generalizability of current approaches.
+Quantifying this imbalance shows that although research has advanced techniques (from heuristic to Transformer models), it has largely done so within a narrow scope of artifacts.
+Highlighting this gap provides guidance for future work to expand the examination of underutilized artifacts, thereby enabling a more comprehensive understanding of SATD.
+Answer to RQ1 Over time, SATD detection has evolved from heuristic methods to ML and DL, with Transformer now leading due to its strength in capturing complex textual dependencies.
+The use of multiple artifacts has increased, yet most studies still focus on code comments, leaving issues, pull requests, and commit messages underrepresented.
+Quantifying this imbalance underscores the need for future research to broaden its focus toward diverse artifacts.
 4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
 To address this, we categorized the methodologies into three main types: manual, semi-automated, and automated.
-These categories represent the evolution in improving SATD detection’s accuracy, efficiency, and scalability.
+These categories represent the evolution in improving the accuracy, efficiency, and scalability of SATD detection.
 Manual methods depend on human expertise to review text artifacts, while semi-automated techniques combine human input with tools to streamline the process.
 Fully automated methods use advanced algorithms to detect SATD with minimal human intervention.
-Each approach has its own strengths and challenges, which we will explore.
-Disagreements in categorization were resolved through discussion.
-Figure 4: Articles count throughout the years based on the methodologies
-Over the past decade, methods for detecting Self-Admitted Technical Debt (SATD) have changed significantly, as shown in Figure 4.
+Each approach has its strengths and challenges, which we will explore.
+To ensure the reliability of this categorization, we followed the procedures described in Subsection 3.6, which included test-retest and cross-checking by the second author.
+No disagreements were recorded during this process, indicating that the grouping was consistent and reliable.
+Figure 4: Distribution of articles by methodology over the years.
+Over the past decade, methods for detecting SATD have changed significantly, as shown in Figure 4.
 There has been a clear shift from manual to more automated techniques, driven by a growing need for efficiency and accuracy in managing technical debt in software development.
@@
 This process involved developers or analysts closely examining codebases and documentation to spot technical debt.
-While these manual methods provided valuable insights, they were time-consuming and susceptible to human error, highlighting the demand for more efficient solutions.
+While these manual methods provided valuable insights, they were time-consuming and prone to human error, underscoring the need for more efficient solutions.
 Between 2017 and 2019, the field began transitioning to semi-automated techniques.
@@
 The figure illustrates a gradual increase in the adoption of semi-automated approaches during this time, often utilizing regular expressions or Part-of-Speech (POS) tagging to reduce developer workload and improve consistency.
-The period from 2020 to 2022 marked a significant move towards fully automated SATD detection methods.
-As depicted in the figure, there was a notable rise in the use of automated techniques, indicating the maturation of the field.
-These approaches employ advanced technologies such as natural language processing (NLP), machine learning (ML), and deep learning (DL) to automatically detect technical debt with minimal human input.
-As we move into 2023 and 2024, the trend toward automation in SATD detection continues to grow.
-The figure shows that automated methods have become the primary approach, outpacing manual and semi-automated techniques.
-This suggests that the future of SATD detection will likely focus on refining automated tools, possibly integrating more advanced AI and ML models.
+The period from 2020 to 2022 marked a significant shift toward fully automated SATD detection methods.
+As shown in the figure, there has been a notable increase in the use of automated techniques, reflecting the maturation of the field.
+These approaches employ advanced technologies such as ML and DL to automatically detect technical debt with minimal human intervention.
+From 2023 onward, automated methods have become the dominant approach, clearly surpassing manual and semi-automated techniques.
+In both 2024 and early 2025, automated detection continued to expand, confirming its position as the leading methodology.
+This trend suggests that the future of SATD detection will likely focus on refining automated tools, with greater reliance on AI-driven techniques to enhance accuracy, efficiency, and adaptability.
 As presented in Table 9, the methodologies are classified into three broad categories: Manual, Semi-Automated, and Automated approaches.
-These categories are determined based on the extent of human intervention and the utilization of NLP, ML, DL, Transformers models.
-Table 8: Methodologies identified
-Methodology Paper Reference NLP ML DL TRANSFORMERS
-Manual No. 8, 37, 59, 39, 102 No. 8, 37, 59, 39, 102 - - -
-Semi-Automated No. 36, 38, 40, 61, 41, 44, 45, 65, 46, 47, 49, 48 No. 36, 38, 40, 41, 44, 45, 46, 47, 49, 48 No. 65, 61 - -
+These categories are determined by the extent of human intervention and the use of Heuristic, ML, DL, and Transformer models.
+Table 9: Methodologies identified.
+Methodology Paper Reference Heuristic ML DL Transformer
+Manual No. 8, 25, 27, 49, 111, 60 No. 8, 25, 27, 49, 111, 60 - - -
+Semi-Automated No. 47, 48, 50, 69, 51, 54, 55, 73, 56, 57, 59, 58 No. 47, 48, 50, 51, 54, 55, 56, 57, 59, 58 No. 73, 69 - -
 Automated No.
-50, 51, 52, 53, 54, 55, 56, 57, 77, 58, 79, 60, 62, 42, 43, 63, 64, 80, 81, 82, 83, 74, 84, 101, 85, 86, 87, 66, 88, 67, 89, 73, 68, 69, 76, 90, 91, 92, 70, 93, 94, 75, 95, 96, 99, 100, 71, 72, 97, 98 - No.
-50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 43, 64, 84, 101, 66, 67, 73, 68, 69, 70, 71, 72 No.
-77, 78, 79, 62, 63, 80, 81, 82, 83, 74, 85, 86, 87, 88, 89, 76, 90, 91, 92, 94, 96 No.
-93, 75, 95, 99, 100, 97, 98
-Manual approaches significantly depend on human intervention, usually requiring the manual annotation of code comments or other textual software artifacts to detect SATD.
-Manual methods solely rely on NLP techniques.
-There is no application of ML, DL, or Transformers in the manual category, as these techniques are generally used for more automated processes.
+26, 24, 61, 62, 63, 64, 65, 66, 85, 67, 87, 68, 70, 52, 53, 71, 72, 88, 89, 90, 91, 83, 92, 110, 93, 94, 95, 28, 96, 74, 97, 82, 75, 76, 30, 98, 99, 100, 77, 101, 102, 84, 104, 105, 108, 10, 78, 79, 106, 107, 31, 80, 103, 81, 109 - No.
+26, 24, 61, 62, 63, 64, 65, 66, 67, 68, 69, 52, 53, 72, 92, 110, 28, 74, 82, 75, 76, 77, 78, 79, 80, 103, 81 No.
+85, 86, 87, 70, 71, 88, 89, 90, 91, 83, 93, 94, 95, 96, 97, 30, 98, 99, 100, 102, 105, 31, 103, 81 No.
+101, 84, 104, 108, 10, 106, 107, 31, 109
+Manual approaches significantly depend on human intervention, typically requiring manual annotation of code comments or other textual software artifacts to detect SATD.
+Manual methods rely solely on heuristic-based techniques.
+There is no application of ML, DL, or Transformer in the manual category, as these techniques are generally used for more automated processes.
 These methods are frequently time-consuming but offer high precision in identifying subtle or context-specific technical debt items.
@@
 They manually classified comments as SATD or non-SATD based on specific comment patterns.
-Similarly, Maldonado and Shihab 37 employed manual labeling to construct a dataset for SATD that focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt.
-This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 51.
+Similarly, Maldonado and Shihab 25 employed manual labeling to construct an SATD dataset focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt.
+This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 24.
 Semi-automated approaches combine manual review with automated tools to improve efficiency while retaining accuracy.
-Semi-automated methods predominantly rely on NLP, including part-of-speech (POS) tagging 36, and textual patterns 38; and a few of them use ML approaches 65, 61.
-However, there is no indication of the usage of Deep Learning or Transformers in these semi-automated techniques.
-Automated approaches are characterized by the use of ML, DL, or Transformers and advanced NLP techniques to fully automate the SATD detection process.
-In order to detect SATD from issue tracking systems, Dai and Kruchten 50 created an automated method that makes use of an ML algorithm, namely of a Naïve Bayes (NB) classifier in addition to other NLP approaches like TF-IDF and TextRank.
-This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1-score.
+Semi-automated methods predominantly rely on heuristic-based techniques, including part-of-speech (POS) tagging 47, and textual patterns 48.
+While some approaches incorporate ML techniques 73, 69, there is no evidence that DL or Transformer-based models are used in these semi-automated techniques.
+Automated approaches are characterized by the use of ML, DL, or Transformer and advanced NLP techniques to fully automate SATD detection process.
+In order to detect SATD from issue tracking systems, Dai and Kruchten 26 created an automated method that uses an ML algorithm, specifically, a Naïve Bayes (NB) classifier, along with other NLP approaches such as TF-IDF and TextRank.
+This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1 score.
 Recently, Li et al.
 30 introduced a novel approach to automatically detect SATD from four different sources: source code comments, commit messages, pull requests, and issue tracking systems.
-The most recent trend in automated SATD detection includes the use of Transformers-based models (e.g., BERT 93, 75, RoBERTa 98, 95, and GPT 99).
-Table 9: Textual artifacts Used. Studies that use more than one textual software artifact are highlighted with square brackets
-Artifact Paper Reference NLP ML DL TRANSFORMERS
-Code comment No.
-8, 36, 37, 38, 51, 52, 53, 54, 55, 56, 57, 77, 58, 78, 40, 40, 60, 61, 62, 41, 42, 63, 64, 80, 81, 44, 82, 83, 45, 84, 101, 85, 86, 65, 66, 88, 67, 89, 73, 46, 68, 91, 92, 70, 93, 94, 75, 96, 99, 71, 49, 48, 72, 97,[69, 76, 90, 100] 8, 36, 37, 38, 40, 41, 44, 45, 46, 49, 48 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 64, 84, 101, 65, 66, 67, 73, 68,[69], 70, 71, 72 77, 78, 79, 62, 63, 80, 81, 82, 83, 86, 88, 89,[76], [90], 91, 92, 94, 96 93, 75, 99,[100], 97
-Issue No. 50, 59, 39, 43, 102, 74, 87, 95, [69, 76, 90, 100] 59, 39, 102, 45 50, 84,[69] 74, 87,[76, 90] 95,[100]
-Pull request No. 98,[69, 76, 100] - [69] [76] 98,[100]
-Commit message No. 47,[69, 76, 100] 47 [69] [76] [100]
-Moving forward from the methodological categorization, it is essential to understand the source or type of textual artifacts analyzed in the context of SATD.
-Table 9 categorizes the various studies on SATD detection based on the types of textual artifacts utilized.
-Code comments (such as from Java projects, blockchain projects, Dockerfile, and XML files), issues (from issue tracking systems and GitHub issue reports), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed.
-The application of NLP, ML, DL, and Transformers-based models across these artifacts shows how methodologies have evolved to handle different levels of complexity and scale.
-Code comments are the most commonly analyzed artifact for SATD detection, with NLP, ML, DL, and Transformers models being widely employed (e.g., studies 8, 36, 93).
-While issues, pull requests, and commit messages are also explored, they are less frequently studied.
-Early research on SATD primarily relied on NLP to extract patterns from developer language in code comments and issues.
-However, more recent studies have incorporated ML, DL, and Transformers models (e.g., 76, 93, 98) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets.
-The increasing adoption of Transformers models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
-Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD.
-Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets.
-The increased use of pull requests and commit messages (e.g., 69, 76, 100) as sources of SATD detection in recent studies signals a growing recognition of these artifacts’ value in providing temporal and contextual data on technical debt.
-As the field moves forward, there is an increasing emphasis on automation and leveraging diverse textual artifacts, which aims to improve the scalability and reliability of SATD detection across varied software projects.
-The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
+The most recent trend in automated SATD detection includes the use of Transformer-based models (e.g., BERT 101, 84, 31, RoBERTa 107, 104, DistilRoBERTa 109, and ChatGPT 108).
 Answer to RQ2 Over the past decade, there has been a visible shift from manual towards fully automated approaches to improve efficiency and scalability in SATD detection.
-Additionally, the use of multiple textual artifacts has grown, enhancing the accuracy and robustness of SATD detection by integrating diverse sources from the software development process.
-4.3 Results - RQ3: Effectiveness of the proposed approaches regarding their metrics evaluation results
-To evaluate the effectiveness of the proposed SATD detection approaches, we examined the performance metrics reported in the studies included in our review.
-While a subset of the studies did not provide detailed performance results, the majority (49 out of 68) reported on key evaluation metrics such as precision, recall, and F1-score.
-The following section presents a detailed analysis of these metrics, utilizing boxplots to visualize the performance variability and trends across different SATD detection approaches.
-Out of 68 relevant studies, 17 did not report their performance results, and 2 only provided AUC scores.
-Therefore, 49 out of the 68 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1-score.
-We collected all performance results presented in these 49 articles and organized them into a spreadsheet.
+4.3 Results - RQ3: Synthesis of reported effectiveness metrics for SATD detection techniques
+As emphasized in Section 1, a key contribution of this review is the systematic compilation and categorization of reported performance scores across SATD detection techniques.
+Rather than conducting new experiments or proposing a benchmarking study, we provide a structured cross-technique perspective on how detection approaches have been reported to perform in the literature.
+To address RQ3, we synthesized the performance metrics as reported in the selected studies.
+It is important to note that the synthesized results are limited to within-project evaluations, typically performed using k-fold cross-validation or random splits on the same dataset or project.
+While some primary studies also reported cross-project or out-of-distribution evaluations (e.g., training on one project and testing on unseen projects), these were deliberately excluded to ensure comparability across studies.
+Consequently, the reported precision, recall, and F1 scores should be interpreted as reflecting performance under controlled within-project settings, which often yield higher values than cross-project evaluations and may overestimate real-world generalizability.
+Following the approach of prior prominent SLRs 112, 113, 114, we synthesized evidence from diverse data sources and models by compiling results reported in the primary studies, rather than conducting new experiments or direct comparisons.
+Because these studies vary in dataset characteristics, size, and benchmarking protocols, the aggregated results should be interpreted as indicative trends rather than definitive rankings.
+Where possible, we report the range and central tendency of precision, recall, and F1 scores to illustrate how different approaches have performed in their original contexts.
+Out of 74 relevant studies, 17 did not report their performance results, and 3 only provided AUC scores.
+Therefore, 54 out of the 74 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1 score.
+We compiled all performance results reported in these 54 articles and organized them in a spreadsheet.
 Rather than focusing solely on the best-performing model from each study, we included the results from all model classifiers.
-This approach provides a more comprehensive and unbiased evaluation of SATD detection models 103.
-By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness.
-Limiting the analysis to only the best-performing models could lead to skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges.
+This approach provides a more comprehensive and unbiased summary of SATD detection models.
+By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness 112.
+Limiting the analysis to only the best-performing models could lead to a skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges.
 Moreover, analyzing all models enables a more robust cross-study comparison, allowing for the identification of consistent trends and patterns in model performance across different contexts, datasets, and techniques.
-This holistic approach strengthens the reliability of the conclusion and ensures their applicability across a wider range of scenarios, rather than being limited to cases of optimal performance.
-We use boxplots to demonstrate the effectiveness of various approaches, as they are a useful graphical tool for showing variations between populations without relying on assumptions about data distribution 103.
-We compiled the effectiveness of various techniques as reported in the original papers, utilizing precision, recall, and F1-score derived from calculations based on the confusion matrix.
-Through boxplots, we showcase studies on precision, recall, and F1-score in relation to different factors of the models.
+This holistic approach strengthens the reliability of the conclusion and ensures its applicability across a broader range of scenarios, rather than being limited to cases of optimal performance.
+We used boxplots to visualize the synthesized results because they are useful for showing variation across populations without assuming specific data distributions 112.
+We extracted the reported effectiveness of each technique from the primary studies, focusing on precision, recall, and F1 scores computed from the confusion matrix.
+The boxplots present the distributions of these metrics across studies and enable comparisons by model-related factors.
   • Precision = $\frac{TP}{TP+FP}$
   • Recall = $\frac{TP}{TP+FN}$
-  • F1-score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
-We acknowledge the potential for misinterpretation when aggregating the results of approaches across various studies without considering the diversity of datasets utilized.
-The rationale behind this methodological choice is to cultivate an inclusive understanding of each model’s performance and versatility across a wide range of contexts and scenarios.
+  • F1 score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
+We acknowledge the potential for misinterpretation when aggregating the results of techniques across various studies without considering the diversity of datasets used.
+We chose this approach to provide a more comprehensive understanding of each model’s performance and versatility across a wide range of contexts and scenarios.
 This inclusive approach enables the identification of universally effective strategies as well as those that excel under specific conditions, thereby providing a nuanced perspective on the potential applicability, hereby offering a holistic view of their effectiveness in detecting SATD.
-This methodology aligns with precedents in the field, notably reflected in the systematic literature review, such as the study of Hall et al.
-103 and Malhotra 104.
+This methodology aligns with precedents in the field, as reflected in systematic literature reviews such as those by Hall et al.
+112 and Malhotra 113.
 This precedent underscores the validity and relevance of cross-dataset comparisons in extracting meaningful insights that transcend the particularities of individual studies.
-By synthesizing findings across diverse datasets, this review aspires to contribute to the advancement of SATD detection approaches, offering valuable guidance for future research and practical applications in software engineering.
-To avoid any misleading interpretations, we divide the process of SATD identification into two distinct tasks: SATD identification and SATD categorization.
-SATD identification focuses on determining whether a particular entity or instance is present in the text of a software development artifact, classifying it as either SATD or non-SATD.
+By synthesizing findings across diverse datasets, this review aims to advance SATD detection techniques and provide valuable guidance for future research and practical applications in software engineering.
+To avoid any misleading interpretations, we divide the process of SATD detection into two distinct tasks: SATD identification and SATD categorization.
+SATD identification aims to determine whether a particular entity or instance is present in the text of a software development artifact and to classify it as SATD or non-SATD.
 SATD categorization, on the other hand, takes this a step further by not only identifying the presence of SATD but also classifying it into specific types of technical debt.
-These categories may include labels such as code debt, design debt, requirement debt, and others, based on the linguistic elements found in the text.
+These categories may include labels such as design debt, architecture debt, and requirement debt, based on the linguistic elements identified in the text.
 In the following subsection, we present performance results for each SATD detection task, illustrated through boxplots.
@@
 The boxplot shown in Figure 5 provides a comparative overview of performance across different SATD identification methods, specifically for binary classification, which aims to accurately distinguish between SATD and non-SATD instances.
-The three key metrics displayed in the box plot—precision, recall, and F1-score—are standard measures for evaluating the effectiveness of classification models.
-Precision indicates the model’s ability to correctly identify instances as SATD or non-SATD, while recall measures how effectively the model captures all relevant SATD instances.
-The F1-score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other.
+The three key metrics displayed in the boxplot, namely precision, recall, and F1 score, are standard measures for evaluating the effectiveness of classification models.
+Precision indicates the model’s ability to correctly identify instances as SATD, whereas recall measures how effectively the model captures all relevant SATD instances.
+The F1 score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other.
 Together, these metrics provide valuable insights into the performance of SATD detection methodologies in identifying technical debt within software projects.
-Figure 5: Boxplot of SATD identification grouped by their approaches
-The NLP-based approach demonstrates a relatively high precision of 0.79, indicating that it successfully identifies SATD in a significant number of cases when it classifies an instance as SATD.
-However, its recall is lower at 0.62, suggesting that it overlooks a considerable portion of SATD instances.
-The F1-score of 0.66 reflects the imbalance between precision and recall, highlighting that while NLP methods were foundational in early SATD detection, they now face challenges in fully capturing all relevant SATD instances.
-In contrast, Machine Learning (ML) models offer a more balanced performance, achieving a precision of 0.73 and recall of 0.74, which results in an F1-score of 0.73.
-This improvement over NLP methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances.
-Historically, ML approaches have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
-Deep Learning (DL) models further enhance performance, with precision and recall scores of 0.78 and 0.76, respectively, yielding a balanced F1-score of 0.76.
-This suggests that DL techniques are better at identifying SATD instances while minimizing false positives.
-The improvements in both precision and recall demonstrate the advancements deep learning architectures have brought to the complexities of SATD identification, enabling more accurate classification across diverse software contexts.
-The Transformers-based approach achieves the best overall performance, with a precision of 0.83—the highest among all methods—and a recall of 0.76, leading to an F1-score of 0.76.
-These results underscore the superior capability of Transformers models to accurately identify SATD while retrieving a majority of relevant instances.
-As the latest advancement in SATD detection techniques, Transformers represent the current state-of-the-art, providing both the precision needed to avoid false positives and the recall required to capture a wide range of SATD instances.
+Finally, we also report the number of experiments (n) for each technique, giving greater weight to algorithms evaluated more frequently, since larger n values provide more reliable evidence and more stable performance estimates across trials.
+Figure 5: Boxplot of SATD identification grouped by their techniques.
+The heuristic-based techniques achieve a relatively high precision of 0.72, indicating their ability to correctly identify SATD when classifying an instance as SATD.
+However, their recall is lower at 0.63, suggesting that they overlook a considerable portion of SATD instances.
+The F1 score of 0.58 reflects this imbalance, indicating that although heuristic-based methods were foundational in early SATD detection, they now struggle to capture all relevant SATD instances.
+In contrast, ML models exhibit more balanced performance, achieving a precision of 0.73 and a recall of 0.73, yielding an F1 score of 0.60.
+This improvement over heuristic-based methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances.
+Historically, ML techniques have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
+DL models further improve performance, achieving a precision of 0.78 and a recall of 0.77.
+This yields a balanced F1 score of 0.76.
+These results suggest that DL techniques are more effective at identifying SATD instances while minimizing false positives.
+Improvements in both precision and recall demonstrate how DL architectures have advanced SATD identification by enabling more accurate classification across diverse software contexts.
+Transformer-based techniques achieved the best overall performance with a precision of 0.76, the highest recall of 0.87, and an F1 score of 0.78.
+These results highlight the superior ability of Transformer models to identify most SATD instances while maintaining competitive precision.
+As the latest advancement in SATD detection, Transformers strike a strong balance between minimizing false positives and maximizing the detection of relevant SATD cases.
 Referring to Figure 5, we can further elaborate and illustrate the metric results in more detail, organizing them based on their respective models, as depicted in Figure 6.
-Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
-NLP-based models, particularly Textual Patterns and Annotation Tags, were assessed for their effectiveness.
-Textual Patterns (n=71) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives.
-However, its lower recall (0.62) and F1 score (0.66) suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification.
+Each model is evaluated using a specified number of experiments (denoted by “n=”), thereby highlighting the range of experiments used to measure its performance.
+Heuristic-based techniques, particularly Textual patterns and Annotation tags, were assessed for their effectiveness.
+Textual patterns (n=94) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives.
+However, its lower recall and F1 score suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification.
 In contrast, Annotation Tags (n=21) performed better in recall and F1 score, indicating a more balanced detection capability, although variability was present, as shown by the wider interquartile range.
 ML models exhibited a broader range of performance.
-Among them, Random Forest (RF) (n=34) and Naive Bayes (NB) (n=37) provided the most consistent results, achieving high precision and F1 scores.
-This indicates their effectiveness in reducing false positives while capturing a majority of SATD instances.
-However, models like Support Vector Machine (SVM) (n=13) and k-NN (n=10) showed lower precision and recall, indicating challenges in accurately identifying SATD instances.
-Notably, Auto-sklearn and LightGBM performed well, though their small sample sizes (n=3 and n=1, respectively) limit the generalizability of their results.
-DL models have made significant strides in SATD identification.
-BiLSTM (n=168), CNN (n=151), and LSTM (n=100) achieved high scores across all metrics, demonstrating consistent precision, recall, and F1 scores.
-These models effectively handle complex SATD detection tasks while minimizing false positives.
-However, models like Graph Neural Network (GNN) (n=35), Recurrent Neural Network (RNN) (n=20), and Graph Convolutional Network (GCN) (n=20) exhibited lower performance, particularly in recall, indicating they may miss important SATD instances.
-Models with smaller samples, such as Ladder Network (LN) (n=10) and Gated Graph Neural Network (GGNN) (n=10), also showed variability in performance, likely due to their limited test sizes.
+Among them, Random Forest (RF) (n=23) provided the most consistent results, achieving high precision and F1 scores with relatively narrow variability.
+Logistic Regression (LR) (n=14), while reaching high median values, showed substantial variability across experiments, suggesting less stable performance.
+Maximum Entropy (ME) (n=46) demonstrated moderately consistent outcomes, though with slightly lower scores compared to RF.
+In contrast, models such as Support Vector Machine (SVM) (n=48) and k-NN (n=13) exhibited lower precision and recall, indicating difficulty in accurately identifying SATD instances.
+Naïve Bayes (NB) (n=1) and Sequential Minimal Optimization (SMO) (n=3) were reported only in a few cases, limiting the generalizability of their results.
+Notably, Auto-sklearn (n=27), LightGBM (n=8), and XGBoost (n=9) also achieved strong results, although their relatively small sample sizes make it difficult to draw robust conclusions.
+DL models show heterogeneous performance in SATD identification.
+BiLSTM (n=170) consistently delivers strong medians, with recall and F1 generally higher than precision and with moderate variability.
+CNN (n=212) achieves comparable median scores but exhibits the largest dispersion and many low outliers, indicating less stable results across studies.
+LSTM (n=20) attains solid medians with noticeable spread, while GNN (n=21) also reaches competitive median values but with wide variability and several low-score cases.
+In contrast, RNN (n=100) trends lower overall, and GCN (n=10) is the weakest, particularly in precision and F1.
+Among the stronger DL variants, LN (n=79) and GGNN (n=20) stand out with high median performance; GGNN is especially competitive while remaining reasonably consistent despite its smaller sample size.
 Transformer-based models represent the most recent advancements in SATD identification.
-Among these, BERT (n=69), RoBERTa (n=39), and Transformer (n=40) achieved the highest precision, recall, and F1 scores.
-These models demonstrated remarkable consistency and reliability in identifying SATD instances, with minimal variability, making them leading techniques in SATD detection.
-Other Transformer models, such as ALBERT (n=18), DeBERTa (n=10), and GPT (n=3), also performed well, though slightly less robustly than the BERT-based models.
-DistilBERT (n=2) showed promise, but its limited sample size makes it difficult to draw definitive conclusions about its overall performance.
-Figure 6: Models performance of SATD identification grouped by each approach
+BERT (n=39) and Transformer (n=102) achieved the strongest and most consistent results, with high precision, recall, and F1 scores.
+ALBERT (n=69) achieves generally strong precision and recall, but its F1 scores vary widely, with several low outliers, making its results less consistent than BERT and Transformer.
+RoBERTa (n=15) performed well, though with fewer experiments, while DeBERTa (n=18) showed more variability and somewhat lower scores.
+ChatGPT (n=2) was evaluated in only a very limited number of cases, yielding inconsistent outcomes.
+DistilBERT (n=5) and Flan-T5 (n=3) achieved promising results but cannot be generalized due to their small sample sizes.
+CodeBERT (n=44) and DistilRoBERTa (n=62) demonstrated consistently strong performance, with DistilRoBERTa in particular matching the stability of BERT-based models.
+Figure 6: Model performance of SATD identification grouped by each technique.
 4.3.2 Performance result SATD categorization
-Figure 7 offers insight into the precision, recall, and F1-score of these approaches, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled approaches differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others.
-The metrics provide insights into how effectively these approaches perform in a multiclass classification scenario, where distinguishing between different categories of SATD is more complex than a binary classification.
-Figure 7: Boxplot of precision, recall, and F1-score of SATD categorization
-NLP approaches demonstrate a low Recall of 0.11 and an F1-score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples.
+Figure 7 offers insight into the precision, recall, and F1 score of these techniques, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled techniques differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others.
+The metrics provide insights into how effectively these techniques perform in a multiclass classification scenario, in which distinguishing among different categories of SATD is more complex than in binary classification.
+Figure 7: Boxplot of precision, recall, and F1 score of SATD categorization.
+Heuristic-based techniques demonstrate a relatively high precision of 0.59 but a very low recall of 0.11 and an F1 score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples.
 This limitation affects their effectiveness in real-world applications, particularly in complex multiclass classification tasks.
-In contrast, ML approaches show notable improvement, achieving a Precision of 0.75 and a Recall of 0.57, resulting in a balanced F1-score of 0.52.
-These results suggest that ML models provide better accuracy in SATD classification, although they still miss a considerable number of true positive cases, as indicated by the Recall.
-DL approaches perform well across all metrics, with a strong Precision of 0.77 and a high Recall of 0.76.
-However, the F1-score of 0.62 suggests that while they successfully capture most true positives, there is still potential for improvement in balancing false positives and false negatives.
-Interestingly, Transformer-based approaches, despite being more advanced, exhibit slightly lower overall performance in this instance, with a Precision of 0.69, Recall of 0.59, and an F1-score of 0.54.
-While transformers typically excel in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on their performance in this specific SATD identification task.
-Nonetheless, transformers still outperform NLP approaches and are comparable to ML models in this context.
+In contrast, ML techniques show notable improvement, achieving a precision of 0.68 and a recall of 0.56, resulting in a more balanced F1 score of 0.52.
+These results suggest that ML models achieve higher accuracy in SATD classification than heuristics, although they still miss a considerable number of true positives, as indicated by the recall.
+DL techniques perform well, achieving a precision of 0.70 and a recall of 0.75, yielding an F1 score of 0.59.
+This indicates that DL models capture most true positives (high recall) while maintaining reasonable precision; however, the overall balance (F1) remains moderate-higher than ML (0.52) but slightly below Transformer (0.61).
+Transformer-based techniques, despite being more advanced, achieved a precision of 0.66, a recall of 0.65, and an F1 score of 0.61.
+While their performance is not markedly higher than that of DL in this setting, Transformers demonstrate consistent, balanced results, outperforming heuristic approaches and ML models and providing comparable effectiveness to DL, with a slightly higher F1 score.
+While Transformer typically excels in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on its performance in this specific SATD identification task.
+Nonetheless, Transformer clearly outperforms Heuristic-based techniques, achieves higher overall scores than ML models, and performs on par with DL approaches, with a slightly better F1 score.
 Referring to Figure 7, we can provide a comprehensive evaluation of specific types of SATD categorization, organized by their respective models, as shown in Figure 8.
-Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
-Figure 8: Models performance of SATD categorization grouped by each approach
-In the NLP group, two approaches, Textual patterns (n=46) and Annotation tags (n=2), were evaluated.
-The Textual Patterns approach showed highly variable performance, with precision hovering around 0.6, while the recall and F1 scores dropped significantly, indicating that while this method can accurately categorize SATDs in some cases, it struggles to consistently retrieve all SATD instances.
-The large interquartile range and the presence of outliers highlight the inconsistency of this method in different contexts.
-In contrast, Annotation tags, although evaluated on only two samples, showed relatively stable and high precision, recall, and F1 scores, indicating strong performance for the multiple cases tested, albeit with limited generalizability.
-These findings confirm the limitations of NLP techniques in handling the complexities of SATD categorization.
-While effective in some cases, particularly for high precision, NLP methods often miss critical SATD instances, leading to suboptimal recall and overall lower F1-scores.
-This aligns with the review’s conclusion that NLP approaches, although foundational in early SATD detection, are no longer sufficient for comprehensive and consistent categorization of SATD.
-ML models showed a wide range of results, reflecting the diversity in algorithm complexity and adaptability to SATD categorization.
-Random Forest (RF) (n=114) and Auto-sklearn (n=108) showed the highest performance, consistently providing high precision and recall, indicating that these models are able to balance false positives and missed SATD examples effectively.
-However, Naive Bayes (n=17) showed weaker performance, particularly in terms of recall, indicating that it may miss important SATD cases.
-Other models, such as SVM (n=25), XGBoost (n=50), and Logistic Regression (n=24), showed lower recall and more significant performance variation across the dataset, reflecting their challenges in categorizing SATD consistently.
-DL models showed significant improvement over the NLP and ML approaches, specifically in their ability to maintain high precision, recall, and F1 scores across larger data sets.
-CNN (n=163), CNN-BiLSTM (n=20), and LSTM (n=70) stood out for their consistent and robust performance, with higher recall rates indicating their effectiveness in detecting a wider range of SATD events without significantly increasing false positives.
-GGNN and Feedforward neural network (FFNN), despite their smaller sample sizes, also showed strong performance, although the small dataset makes it difficult to draw broader conclusion about their generalizability.
-Transformers-based models, specifically BERT (n=47), ALBERT (n=24), and RoBERTa (n=13), were the highest performing approaches in SATD categorization.
-BERT emerged as the best performer, with high precision, recall, and F1 scores across all evaluations, demonstrating its ability to accurately and consistently detect SATDs with minimal false positives and missed events.
-ALBERT and RoBERTa also performed strongly, although evaluated on smaller sample sizes, their high precision and F1 scores reinforced their value in SATD detection.
-From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1-score.
+Each model is evaluated using a specified number of experiments (denoted by “n=”), thereby highlighting the range of experiments used to measure its performance.
+Figure 8: Model performance of SATD categorization grouped by each technique.
+Heuristic-based techniques were evaluated using Textual patterns (n=46) and Annotation tags (n=6).
+Textual patterns showed highly variable performance, with precision around 0.6 but much lower recall and F1 scores, indicating that while this method can sometimes classify SATDs accurately, it struggles to consistently retrieve all instances.
+The wide interquartile range and presence of outliers highlight its inconsistency across contexts.
+In contrast, Annotation tags, although tested in only six cases, showed relatively stable performance, with higher precision, recall, and F1 scores, suggesting stronger performance but limited generalizability.
+These findings confirm the limitations of heuristic approaches: while sometimes effective for achieving high precision, they often miss critical SATD instances, leading to suboptimal recall and lower overall F1 scores.
+ML models exhibited a broad spectrum of results, reflecting their algorithmic diversity.
+Random Forest (RF) (n=114) and Auto-sklearn (n=114) achieved the strongest performance, with consistently high precision and recall, effectively balancing false positives and missed SATD cases.
+In contrast, Naïve Bayes (n=17) performed more weakly, especially in recall.
+Other models, such as Support Vector Machine (SVM) (n=25), Logistic Regression (LR) (n=24), and XGBoost (n=58), showed lower recall and greater variability, indicating challenges in reliably categorizing SATD.
+Maximum Entropy (ME) (n=18) demonstrated moderate performance, whereas Sequential Minimal Optimization (SMO) (n=8) was tested in only a limited number of cases, thereby limiting its generalizability.
+DL models demonstrated clear improvements over heuristic and ML techniques.
+CNN (n=205) achieved strong and consistent results across metrics, while LSTM (n=70) also performed well but with greater variability.
+Gated Graph Neural Network (GGNN) (n=20) showed promising outcomes, though based on fewer samples.
+Feedforward Neural Network (FFNN) (n=4) displayed relatively strong performance, but the extremely small sample size limits conclusions about its broader applicability.
+Transformer-based models emerged as the most advanced group.
+BERT (n=28) achieved consistently high precision, recall, and F1 scores, standing out as the top performer.
+ALBERT (n=23) also performed strongly, though with more variability, while RoBERTa (n=13) maintained high scores despite its smaller sample size.
+Additional models, including CodeBERT (n=43) and Flan-T5 (n=22), also demonstrated solid and balanced performance, reinforcing their reliability for SATD detection.
+DistilRoBERTa (n=2), however, was evaluated on very limited data, making its results less conclusive.
+From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1 score.
 It is important to note that Figure 6 represents SATD identification aimed at distinguishing SATD from non-SATD (binary classification), while Figure 8 illustrates the results of SATD categorization of specific SATD types (multiclass classification).
-To select the best-performing algorithm from the boxplot, we primarily focused on the F1-score (dark blue bars) as it offers a balance between precision and recall.
-We identified the algorithm with the highest median F1-score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance suggesting more consistent performance.
-Additionally, we examined the precision and recall (light and mid-blue bars) to ensure a balance between the two, prioritizing algorithms with high values in both.
-In addition, we minimized the influence of algorithms with numerous outliers, as they indicate inconsistent performance, even if their median scores were high.
-Finally, we factored in the n value (number of experiments), giving preference to algorithms tested more frequently, as a higher n suggests more reliable and stable results across different trials.
-Table 10: Summary of highest performance results based on SATD detection tasks
-Approach Metric Evaluations
-Precision Recall F1-score
+To select the best-performing algorithm from the boxplot, we primarily focused on the F1 score (dark blue bars) as it offers a balance between precision and recall.
+We identified the algorithm with the highest median F1 score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance indicating more consistent performance.
+Additionally, we examined precision and recall (light and mid-blue bars) to ensure a balanced trade-off between the two, prioritizing algorithms with high values for both.
+In addition, we minimized the influence of algorithms with numerous outliers, as such algorithms indicate inconsistent performance, even when their median scores are high.
+Table 10: Summary of highest performance results based on SATD detection tasks.
+Technique Metric Evaluations
+Precision Recall F1 score
 SATD identification
-NLP Annotation tags Annotation tags Annotation tags
-ML SMO LR SMO
-DL GNN BiLSTM BiLSTM
-Transformers GPT Transformer BERT
+Heuristic Annotation tags Annotation tags Annotation tags
+ML RF RF RF
+DL CNN BiLSTM BiLSTM
+Transformer DistilRoBERTa DistilRoBERTa DistilRoBERTa
 SATD categorization
-NLP Annotation tags Annotation tags Annotation tags
-ML SMO NB SMO
+Heuristic Annotation tags Annotation tags Annotation tags
+ML Auto-sklearn Auto-sklearn Auto-sklearn
 DL GGNN GGNN GGNN
-Transformers BERT BERT BERT
-The best-performing algorithms for each approach based on the different SATD types is presented in Table 10.
-In both the SATD identification and categorization tasks, BERT consistently emerges as the best-performing model.
-For SATD identification, BERT’s high median F1-score, tested across a large number of experiments (n=69), showcases its strong reliability and generalizability.
-Its low variance and minimal outliers further demonstrate its consistent and stable performance across different datasets.
-Leveraging its advanced Transformer-based architecture, BERT excels at capturing the context and nuances in SATD instances, making it highly effective in balancing precision and recall.
-Similarly, in the SATD categorization task, BERT outperformed other models, achieving the high median F1-score with a significant number of experiments (n=47).
-Its minimal variance and absence of outliers reinforce its reliability and consistency in this task as well.
-Altogether, these factors position BERT as the most robust and effective model for both SATD identification and categorization, making it the top choice in each task.
-Answer to RQ3 BERT (a Transformers-based model) emerges as the top performer in both SATD and non-SATD identification and the categorization of specific SATD types, providing balanced precision, recall, and F1 scores.
-SMO (an ML model) and GGNN (a DL model) also show strong performance within their respective domains, with SMO excelling in precision and F1 score, while GGNN leads in identifying specific SATD types.
-Overall, Transformer models, particularly BERT, represent the most advanced and effective approach to SATD detection.
+Transformer BERT BERT BERT
+The best-performing algorithms for each technique and SATD task are summarized in Table 10.
+Overall, Transformer-based models consistently deliver the strongest reported results across both SATD identification and categorization.
+For SATD identification, DistilRoBERTa yields the highest precision, recall, and F1 score among Transformer variants in our synthesis, indicating strong and balanced effectiveness in binary classification settings.
+ML models, particularly Random Forest, also show competitive performance for identification, especially in precision and F1, while among DL models BiLSTM achieves the strongest overall balance between precision and recall.
+For SATD categorization, BERT remains the best-performing Transformer model, achieving the highest precision, recall, and F1 score, while Auto-sklearn and GGNN lead the ML and DL groups, respectively.
+These outcomes reflect trends reported across the literature rather than a definitive ranking, as performance is strongly influenced by dataset quality, annotation practices, and evaluation protocols.
+Answer to RQ3 Overall, Transformer-based models achieve the strongest reported effectiveness across SATD detection tasks.
+In SATD identification, DistilRoBERTa attains the highest precision, recall, and F1 score in the synthesized results, while in SATD categorization BERT is the best-performing model across all three metrics.
+Although several ML (e.g., RF and Auto-sklearn) and DL models (e.g., BiLSTM and GGNN) also achieve strong results in specific settings, their performance is generally less consistent than Transformer-based techniques across the heterogeneous datasets and evaluation protocols used in the primary studies.
 4.4 Results - RQ4: Proposed tools for SATD detection
-To address this research question, we conducted a thorough review of existing tools developed over the past decade.
-While the detection and management of SATD have been extensively studied, there remains a noticeable gap in translating research into practical tools for developers.
-This section examines tools specifically designed for SATD detection, offering insights into advancements in automation and machine learning, as well as the diverse approaches used to address various forms of technical debt.
-Table 11 presents a collection of tools proposed for SATD identification across different software projects.
-This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, machine learning integration, and varied strategies for managing different types of technical debt.
-Below, we describe these tools and their significance in the broader context of SATD identification, emphasizing key trends that have emerged over the last decade.
-Table 11: SATD detection tools proposed
-Name Reference Year Task Type Description Status
-eXcomment No. 36, 48 2015 Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
-SATD Detector No.
-55 2018 Identification A Java library and an Eclipse plug-in.
+To address this research question, we conducted a comprehensive review of existing tools developed over the past decade.
+While SATD detection and management have been extensively studied, there remains a significant gap in translating research findings into practical, developer-friendly tools.
+This section focuses on tools specifically designed for SATD detection, providing insights into advancements in automation and the various techniques used to address different forms of technical debt.
+In this SLR, we classify an approach as a tool only when the authors explicitly propose a prototype or a ready-to-use system, rather than merely providing source code.
+To be considered a valid contribution to SATD detection, we regard a tool as a system designed for external use, meaning it can be executed with minimal setup and without requiring significant modifications and must be deployable and practically usable 115, 116.
+While some studies provide source code to support their experiments, the availability of source code alone is not sufficient to qualify an approach as a tool.
+Additionally, we include studies in which a tool is described and a replication package is provided, even if the authors do not explicitly refer to it as a tool or if the provided link is no longer valid, as long as the paper provides sufficient evidence that the tool was originally intended for SATD detection.
+Table 11 presents a collection of SATD detection tools proposed in different software projects.
+This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, ML integration, and varied strategies for managing different types of technical debt.
+The following discussion examines the significance of these tools within the broader context of SATD detection and identifies key trends that have shaped their development over the past decade.
+Table 11: SATD detection tools proposed.
+Name Year Technique Automation Level Artifact Task Type Description Status
+BEACon-TD 109 2025 Transformer Automated Issue Identification, Categorization DistilRoBERTa-based tool that classifies software issues as technical debt and further categorizes them into 13 types using a two-stage ensemble of binary classifiers Available
+DebtViz 98 2023 DL Automated Code comment & Issue Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
+SATDBailiff 74 2022 ML Automated Code comment Identification A tool designed to mine, identify, and track SATD Available
+A browser extension 83 2022 ML Automated Issue Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of an rOpenSci R package Not available
+DebtHunter 72 2021 ML Automated Code comment Identification, Categorization A machine learning-based technique for detecting self-admitted technical debt Available
+FixMe 79 2021 ML Automated Code comment Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
+SATD Detector 64 2018 ML Automated Code comment Identification A Java library and an Eclipse plug-in.
 The tool automatically identifies whether comments are SATD or non-SATD and integrates with an IDE to help manage these comments Obsolete
-DebtHunter No. 64 2021 Identification, Categorization A machine learning-based approach for detecting self-admitted technical debt Available
-FixMe No. 72 2021 Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
-SATDBailiff No. 67 2022 Identification A tool designed to mine, identify, and track SATD Available
-A browser extension No.
-74 2022 Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of a rOpenSci R package Not available
-DebtViz No.
-90 2023 Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
-The tools proposed for SATD identification reflect a diverse range of approaches to addressing technical debt in software projects.
-eXcomment 36, 48, and SATD Detector 55 focus on identifying both SATD and non-SATD comments, with eXcomment integrating into an IDE for real-time detection.
-A browser extension 74 leverages machine learning to detect various SATD types, including documentation and usability debt, highlighting the potential of ML for broader SATD detection.
-SATDBailiff 67 provides comprehensive tracking and monitoring of SATD over time, while FixMe 72 is a GitHub bot specifically designed to alert developers about unresolved, “on-hold SATD" in repositories.
-Lastly, DebtViz 90 offers a more complete solution, combining SATD detection, classification, visualization, and monitoring in one platform, emphasizing the need for visual tools that make technical debt more understandable to stakeholders.
-Together, these tools demonstrate an increasing focus on automation, monitoring, and comprehensive SATD management, though the overall number of tools remains limited relative to the research interest in the field.
-This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 68 relevant studies, only 7 articles proposed specific tools for SATD identification.
-This disparity points to a concerning gap between the significant research interest in SATD and the actual development of practical, automated solutions that can assist developers in managing technical debt.
-While a range of tools has been introduced, the overall lack of available tools indicates that much of the research has yet to translate into actionable, industry-ready solutions.
+eXcomment 47, 58 2015 Heuristic Semi-Automated Code comment Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
+The proposed tools for SATD identification employ diverse techniques and varying levels of automation to address technical debt in software projects.
+eXcomment 47, 58 utilizes a heuristic approach with semi-automated functionality to parse Java source code comments and detect SATD.
+In contrast, SATD Detector 64 applies ML in a fully automated manner to identify and classify SATD derived from code comments within an IDE, although it has since become obsolete.
+Advancing the role of ML, tools such as DebtHunter 72 and FixMe 79 implement automated SATD detection and categorization.
+DebtHunter employs ML to detect SATD, while FixMe functions as a GitHub bot that actively monitors and notifies developers about unresolved On-hold SATD.
+SATDBailiff 74 further extends SATD management by introducing an automated tool designed to mine, track, and monitor SATD occurrences over time.
+Another approach is demonstrated by a browser extension 83, which leverages an ML model to automatically categorize different SATD types, such as documentation, usability, and requirement debt.
+Lastly, DebtViz 98 adopts DL to provide an automated solution that integrates detection, classification, visualization, and monitoring of various SATD types within a single platform.
+More recently, BEACon-TD 109 introduces a Transformer-based solution that detects SATD and classifies it into 13 specific types using a two-stage ensemble of binary classifiers.
+This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 74 relevant studies, only 8 articles proposed specific tools for SATD identification.
+This disparity highlights a concerning gap between the substantial research interest in SATD and the development of practical, automated solutions to assist developers in managing technical debt.
+While a range of tools has been introduced, their limited availability indicates that much of the research has yet to translate into actionable, industry-ready solutions.
 Addressing this gap will be critical for advancing the practical management of technical debt in software projects, moving from theoretical frameworks toward robust, scalable automation.
-Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, machine learning integration, and addressing different types of technical debt.
+Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, ML integration, and addressing different types of technical debt.
 However, despite significant research interest, only a few tools have been developed, revealing a gap between theoretical advancements and the availability of practical, automated solutions for managing technical debt in real-world software projects.
@@
 These challenges reflect the complexities inherent in identifying SATD across various software projects and methodologies.
-We identified nine key challenges related to SATD identification extracted from the relevant studies, as summarized in Table 12, with relevant studies emphasizing each challenge.
-These challenges were determined by examining sections commonly located in the ‘Discussion’ or ‘Threat to Validity’ sections of the papers reviewed.
-We extracted all discernible challenges, whether implicitly or explicitly mentioned, associated with SATD detection.
-Each challenge is explained in this section and further elaborated on below.
-Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets
+We identified nine key challenges related to SATD identification from the relevant studies, as summarized in Table 12, and each study emphasizes one of these challenges.
+The extraction process focused on statements reported in the ‘Discussion’ or ‘Threats to Validity’ sections of the reviewed papers.
+All discernible challenges, whether explicitly or implicitly mentioned, were collected for further analysis.
+The grouping of challenges was also performed using a bottom-up thematic analysis.
+Each extracted challenge statement was coded and then clustered into higher-level categories based on thematic similarity, following common practices in prior SE literature reviews.
+The first author conducted the initial coding and grouping, while the second author independently cross-checked a 15% random sample of the classifications using the shared spreadsheet.
+Any disagreements were resolved through collaborative discussion until a consensus was reached.
+Only one case required clarification, indicating that disagreements were minimal and consistently resolvable.
+Each resulting challenge category is explained in this section and further elaborated on below.
+Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets.
 No Challenge Paper Reference
-1 Error-proneness dues to manual analysis No. 8, 38,[37, 36, 50, 53, 59, 39, 49]
-2 Imbalanced data No. [54, 77, 78, 40, 79, 60, 61, 62, 64, 81, 82, 83, 45, 85, 86, 88, 68, 93, 94, 96, 71], 56, 63, 80, 66, 89, 46, 47, 92, 75, 95, 97
-3 Generalizability No. [50, 53, 59, 39, 37, 55, 77, 40, 60, 61, 64, 78, 81, 45, 85, 86, 88, 68, 93, 94, 96, 49, 99, 71], 58, 74, 101, 87, 65, 69
+1 Error-proneness No. 8, 48,[25, 47, 26, 62, 27, 49, 59]
+2 Imbalanced data No. [63, 85, 86, 50, 87, 68, 69, 70, 72, 89, 90, 91, 55, 93, 94, 96, 75, 101, 102, 105, 78, 65, 71, 88, 28, 97, 56, 57, 100], 84, 104, 106, 80, 60
+3 Generalizability No. [26, 62, 27, 49, 25, 64, 85, 50, 68, 69, 72, 86, 89, 55, 93, 94, 96, 75, 101, 102, 105, 59, 108, 78], 67, 83, 110, 95, 73, 76, 31, 81, 109
 4 Explainability No. [85, 86, 91, 93, 94, 96, 108, 54, 10]
@@
 6 Risk of reproducibility No. [ 54, 10, 74]
-7 Computational costs No. [44, 82, 83, 67, 100]
+7 Computational costs No. [54, 90, 91, 74, 10], 103
 8 Overlap of features among debt types No. [25, 28]
 9 Tagging and comment relevance No. [47, 59]
-Manual Analysis
-The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 38.
-Since the process relies on human judgment, it becomes prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD differently 36.
-This subjectivity in analysis can lead to varying outcomes in tasks like pattern recognition and feature engineering, making the process less reliable.
-The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can severely impact the effectiveness of machine learning models that rely on this data for training.
+Error proneness
+The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 48.
+Since the process relies on human judgment, it is prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD in different ways 47.
+This subjectivity in analysis can lead to varying outcomes in tasks such as pattern recognition and feature engineering, thereby reducing the reliability of the process.
+The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can significantly affect the effectiveness of ML models that rely on this data for training.
 When models are trained on datasets that contain inaccuracies or inconsistencies, their predictive accuracy and generalization capabilities suffer.
 This challenge becomes even more pronounced when scaling up to larger datasets, where consistent and precise labeling is critical to maintaining the quality of the training data.
-Without robust and reliable analysis, the models may struggle to handle the increased complexity of larger data, leading to reduced performance and misidentification of SATD instances.
+Without robust and reliable analysis, models may struggle to handle the increased complexity of large datasets, resulting in reduced performance and misclassification of SATD instances.
 Imbalanced data
 In most software projects, SATD comments represent only a small portion of the total code comments, resulting in a significant class imbalance.
-This imbalance can adversely affect the performance of SATD identification models, as the models may tend to focus on non-SATD comments, which dominate the dataset.
-As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 66.
-This imbalance frequently leads to a high number of false negatives 105, where the model fails to detect SATD instances, undermining its effectiveness.
-Since machine learning models generally find it difficult to learn patterns from underrepresented classes 106, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics.
+This imbalance can adversely affect the performance of SATD identification models, as they may focus on non-SATD comments that dominate the dataset.
+As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 28.
+This imbalance frequently leads to a high number of false negatives 117, in which the model fails to detect SATD instances, thereby undermining its effectiveness.
+Since ML models generally find it difficult to learn patterns from underrepresented classes 118, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics.
 Consequently, the model’s ability to identify and classify SATD instances diminishes, particularly when faced with a large volume of non-SATD items.
-Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that the models are not biased toward detecting the majority class while neglecting the crucial SATD instances.
+Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that models are not biased toward detecting the majority class while neglecting crucial SATD instances.
 Generalizability
-One of the primary challenges associated with the identification of SATD is the issue of generalizability.
-Generalizability refers to the ability of SATD identification models to maintain their accuracy and effectiveness when applied across different software projects and domains.
-This challenge arises due to the diverse nature of software projects, which often exhibit significant variations in coding practices, domain-specific terminologies, and architectural designs 44.
+One of the primary challenges in identifying SATD is its generalizability.
+Generalizability refers to the extent to which SATD identification models maintain their accuracy and effectiveness when applied across different software projects and domains.
+This challenge arises from the diverse nature of software projects, which often exhibit significant variation in coding practices, domain-specific terminology, and architectural designs 54.
 Models trained on data from one specific project tend to be tailored to the characteristics of that project, limiting their performance when applied to other projects with different practices.
-For example, the way SATD is expressed in comments may differ across projects due to differences in the development team’s language, domain jargon, or code style conventions 58, 39.
-These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when faced with unfamiliar coding practices or new domains.
+For example, the way SATD is expressed in comments may vary across projects due to differences in the development team’s language, domain-specific jargon, or coding style conventions 67, 49.
+These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when encountering unfamiliar coding practices or new domains.
 Explainability
-Another primary challenge in the SATD identification process, particularly in machine learning and deep learning-based models, is the lack of explainability.
-This issue arises because many advanced models, such as neural networks and deep learning architectures, function as “black boxes," meaning their internal decision-making processes are not easily interpretable by users 107.
+Another primary challenge in the SATD identification process, particularly in ML and DL-based models, is the lack of explainability.
+This issue arises because many advanced models, such as neural networks and DL architectures, function as “black boxes,” meaning their internal decision-making processes are not easily interpretable by users 119.
 These models generate predictions without providing transparent explanations about the underlying reasoning for classifying an instance as SATD, non-SATD, or even specific types of SATD.
@@
 False negatives are another concern in SATD detection.
-They occur when SATD items are incorrectly classified as non-SATD, leading to undetected technical debt in the codebase.
-This happens when models fail to identify instances where developers have either implicitly or explicitly admitted technical debt 38.
-As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 12, 108.
+They occur when SATD items are misclassified as non-SATD, resulting in undetected technical debt in the codebase.
+This occurs when models fail to identify instances in which developers have either implicitly or explicitly admitted technical debt 48.
+As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 13, 120.
 Risk of reproducibility
-Another primary challenge in the context of machine learning and supervised models-based models is the issue of reproducibility 44, 67.
+Another primary challenge in the context of machine learning and supervised model-based models is the issue of reproducibility 54, 74.
 These models often involve complex architectures with numerous parameters that must be fine-tuned for optimal performance.
 However, many studies fail to provide sufficient details on the hyperparameters, training configurations, or the specific processes used to achieve their results.
-Additionally, the lack of publicly available code exacerbates the reproducibility problem, as other researchers are unable to replicate the exact experimental setup.
-This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these approaches.
+Additionally, the lack of publicly available code exacerbates reproducibility issues, as other researchers cannot replicate the exact experimental setup.
+This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these techniques.
 Computational costs
-Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly deep learning approaches.
+Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly DL and Transformer techniques.
 These models require significant resources during training, including high-performance GPUs, large memory capacities, and extended computation time.
-This high computational cost creates several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
-Additionally, the extended computation time required for training deep learning models presents a challenge in terms of practicality.
-SATD identification models often need to be retrained or fine-tuned to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices.
-In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of deep learning models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
+This high computational cost poses several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
+Additionally, the extended computation time required for training DL models presents a challenge in terms of practicality.
+SATD identification models often require retraining or fine-tuning to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices.
+In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of DL models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
 Overlap of features among debt types
-Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 66.
+Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 28.
 This overlap makes distinguishing between different debt types difficult, leading to potential misclassification.
 The ambiguity in distinguishing between SATD and other debt types can reduce the accuracy of identification models.
-If features that are common to multiple debt types are not carefully handled, models may struggle to correctly identify SATD.
+If features that are common to multiple debt types are not carefully handled, models may struggle to identify SATD correctly.
 Tagging and comment relevance
-Finally, tagging and comment relevance is a critical challenge in identifying SATD.
-Developers often use tags such as “TODO, “FIXME," or “XXX" in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes.
+Finally, tagging and comment relevance are a critical challenge in identifying SATD.
+Developers often use tags such as “TODO,” “FIXME,” or “XXX” in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes.
 While these task-related tags are commonly employed as indicators of SATD, they are not always relevant to technical debt.
-In many cases, tags like “TODO" or “FIXME" can represent a simple reminder or a non-technical debt issue, thereby creating noise in the identification process.
-Moreover, not all SATD instances are explicitly marked by such tags, further complicating the identification process.
-Answer to RQ5 We identified nine major challenges in detecting SATD, including issues like manual analysis errors, imbalanced data, lack of generalizability, and explainability in advanced models.
+In many cases, tags such as “TODO” or “FIXME” can serve as simple reminders or non-technical debt items, thereby creating noise in the identification process.
+Moreover, not all SATD instances are explicitly marked with such tags, further complicating the identification process.
+The results highlight significant trends in SATD detection, demonstrating an evolution from heuristic-based methods to advanced machine learning and deep learning techniques.
+However, these findings raise important questions regarding the applicability, generalizability, and practical adoption of these approaches.
+The following section presents a critical discussion of our results, compares them with the existing literature, and outlines their implications for future research and industry practice.
+Answer to RQ5 We identified nine major challenges in detecting SATD, including manual analysis errors, data imbalance, and limited generalizability and explainability in advanced models.
 Other challenges include false negatives, reproducibility risks, computational costs, overlapping features among debt types, and the relevance of tagging and comments, all of which hinder effective SATD identification.
 5 Discussion
-The findings from this systematic review reveal several key trends and challenges in the detection of SATD over the past decade, demonstrating both the evolution and fragmentation within the field.
-SATD detection approaches have progressed significantly, shifting from simple keyword-based NLP methods to more complex ML, DL, and Transformers-based models.
+The findings from this systematic review reveal several key trends and challenges in detecting SATD over the past decade, demonstrating both the evolution and fragmentation within the field.
+SATD detection techniques have progressed significantly, shifting from simple heuristic-based methods to more complex ML, DL, and Transformer-based models.
 While each advancement has brought increased performance and automation, the development of practical and scalable SATD detection tools remains limited.
-5.1 Evolution of SATD detection techniques
-Early SATD detection techniques primarily relied on basic NLP methods, such as keyword matching and textual patterns, to identify technical debt in code comments.
-These approaches, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts.
-From 2017 onward, ML techniques such as Naive Bayes (NB), Support Vector Machines (SVM), and Random Forests became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives.
+5.1 Evolution of SATD detection techniques and artifacts
+Early SATD detection techniques primarily relied on heuristic-based methods, such as keyword matching and textual patterns, to identify technical debt in code comments.
+These techniques, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts.
+From 2017 onward, ML techniques such as NB, SVM, and RF became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives.
 Despite this progress, these methods still faced challenges related to generalizability across different datasets and projects, as well as difficulty in explaining the rationale behind the classification of certain comments as SATD.
-The rise of deep learning methods from 2020 to 2022 marked a significant leap forward in SATD detection.
-Models like Long Short-Term Memory (LSTM) networks and Convolutional Neural Networks (CNN) enabled the identification of more complex patterns in developer comments and project artifacts.
-These approaches improved recall and reduced false negatives, offering a more balanced approach to identifying SATD across diverse datasets.
+The rise of DL methods from 2020 to 2022 marked a significant leap forward in SATD detection.
+Models like LSTM and CNN enabled the identification of more complex patterns in developer comments and project artifacts.
+These techniques improved recall and reduced false negatives, thereby providing a more balanced technique for identifying SATD across diverse datasets.
 However, the high computational cost of training and deploying these models remains a major barrier to widespread adoption.
-Most recently, Transformers-based models, such as BERT and RoBERTa, have emerged as the leading approaches for SATD detection.
+Most recently, Transformer-based models, such as BERT and RoBERTa, have emerged as the leading techniques for SATD detection.
 These models have demonstrated superior performance in capturing the context and semantics of developer comments, enabling more accurate and generalizable SATD identification.
-The ability of Transformers to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection.
+The ability of Transformer to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection.
 However, issues related to the explainability of these models and their scalability for industrial use still persist.
+It is also important to note that the evolution of these techniques has occurred primarily within the context of code comments, which remain the dominant artifact for SATD detection.
+While this focus is understandable given the accessibility and ubiquity of comments across software projects, it narrows the diversity and generalizability of existing approaches.
+Other artifacts, such as issue trackers, commit messages, and pull requests, have been comparatively underexplored despite their potential to provide complementary perspectives on technical debt.
+Addressing this imbalance represents a valuable direction for future research, as expanding the scope of artifacts can enable a more comprehensive understanding of SATD across the software development process.
 5.2 Challenges in SATD detection
@@
 A key challenge in SATD detection is the imbalance in datasets—SATD instances are often much less frequent than non-SATD instances, which negatively impacts model performance.
-This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, are far less represented, further complicating detection.
+This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, being far less represented, further complicating detection.
 Although several techniques have been employed to address data imbalance in SATD detection, future research should explore more advanced methods, such as leveraging large language models (LLMs) to better handle imbalanced data and improve model performance.
 Generalizability remains another key challenge.
-Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their ability to perform well on new or proprietary codebases.
+Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their performance on new or proprietary codebases.
 This lack of generalizability reduces the practical utility of these models in diverse software development environments, where coding practices, project documentation, and domain-specific language can vary significantly.
-The explainability of advanced models, particularly deep learning and Transformer-based approaches, is also a concern.
-Without a clear understanding of why certain comments are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows.
+To address this, future research should move beyond within-project validation and systematically evaluate SATD detection models “in the wild”, using out-of-distribution datasets and cross-project settings.
+Such evaluations, as demonstrated by recent work like 109, are essential for understanding how well these models transfer to unseen contexts and for ensuring their reliability in real-world adoption.
+The explainability of advanced models, particularly DL and Transformer-based techniques, is also a concern.
+Without a clear understanding of why certain texts are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows.
 Future work should focus on developing explainable AI techniques for SATD detection to ensure that these models’ decision-making processes are transparent and understandable to developers.
-Research on enhancing the interpretability of DL and Transformers models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
-Another significant barrier to adoption is the high computational cost associated with training and maintaining deep learning and Transformers-based models.
+Research on enhancing the interpretability of DL and Transformer models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
+Another significant barrier to adoption is the high computational cost associated with training and maintaining DL and Transformer-based models.
 While these models offer improved accuracy, their resource-intensive nature makes them difficult to implement for smaller development teams or in environments with limited computational resources.
-This challenge highlights the need for more efficient algorithms or alternative approaches that can deliver high accuracy without the associated computational burden.
+This challenge highlights the need for more efficient algorithms or alternative techniques to deliver high accuracy without the associated computational burden.
 5.3 Lack of tools for SATD detection
 Despite the significant academic interest in SATD detection, the translation of research into practical tools remains limited.
-This review identified only seven tools developed specifically for SATD detection over the past decade, reflecting a notable gap between theoretical research and practical application.
-While tools like SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, the overall number of available tools is insufficient to meet the needs of diverse software development teams.
-Additionally, the tools that do exist often lack the ability to integrate seamlessly into modern development environments, limiting their utility for real-time SATD management.
-The limited development of practical tools also highlights the need for further collaboration between academia and industry.
-Researchers must focus on creating tools that not only demonstrate high performance in academic settings but also address the specific needs and constraints of industrial software development, such as scalability, usability, and integration with existing development environments.
-Bridging the gap between theoretical advancements and practical, scalable tools is essential for achieving widespread adoption in industry.
-Future research should prioritize developing user-friendly tools that can be integrated into continuous integration (CI) environments, enabling real-time SATD detection and providing actionable insights for developers.
-Moreover, these tools should facilitate collaboration within teams working on large-scale software projects, ensuring that they are both practical and effective in real-world development scenarios.
+Over the past decade, this systematic review identified only seven tools developed specifically for SATD detection.
+This reflects a notable gap between theoretical research and practical application and a lack of sufficient tools to meet the diverse needs of software development teams.
+While tools such as SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, their scope is limited in terms of flexibility, scalability, and integration with modern software development workflows.
+Consequently, these tools have experienced limited adoption beyond academic prototypes and small-scale studies.
+The current state of SATD detection tools presents several challenges that hinder their widespread adoption among practitioners.
+Key areas of concern include limited real-time integration, inadequate support for multiple programming languages, insufficient collaboration features, and poor scalability for large codebases.
+Modern software development practices are heavily reliant on continuous integration (CI) and continuous deployment (CD) pipelines 121, alongside integrated development environments (IDEs) 122 such as IntelliJ IDEA, Visual Studio Code, and Eclipse.
+However, most SATD detection tools lack seamless integration with these environments, necessitating manual execution or batch processing.
+This disrupts established workflows and reduces the tool’s utility for practitioners who require real-time feedback during development.
+The absence of real-time notifications limits the capacity of developers to address technical debt in a timely manner, thereby contributing to its accumulation.
+Tools designed for practical use should therefore provide IDE integration and support for automated scanning within CI/CD pipelines to enhance workflow efficiency and developer productivity.
+Furthermore, existing tools often struggle to scale effectively in large and complex software projects.
+Many of these tools are designed with a narrow focus on specific programming languages, particularly Java, which limits their applicability in diverse development environments where teams work with multiple languages, such as Python, C++, and JavaScript.
+This restriction reduces their practicality for organizations seeking comprehensive SATD detection across their entire codebase.
+Performance optimization for large repositories and multi-language compatibility are thus critical areas for future improvement.
+In addition, many SATD detection tools rely on ML or DL models, which are often perceived as “black boxes”.
+This lack of interpretability may make it challenging for developers to fully trust and adopt these tools.
+Offering more transparent and explainable outputs—such as highlighted patterns, relevant keywords, or confidence scores—could enhance clarity, build trust, and encourage broader adoption among practitioners.
+Moreover, existing tools seldom offer functionality for long-term monitoring and trend analysis of SATD.
+For projects with extended lifecycles, it is valuable to track how technical debt is introduced, accumulated, and eventually repaid.
+While tools such as DebtViz aim to visualize and monitor SATD over time, they remain in early stages of development and may not yet be suitable for large-scale applications.
+Enhancing these tools with features for historical analysis and prioritization of technical debt repayment could support more informed decision-making and strategic refactoring efforts.
+These limitations highlight a broader challenge: the gap between academic research and industry needs.
+Many tools are developed with a strong focus on academic performance metrics, yet they may not fully address practical constraints, including integration, scalability, and usability.
+Strengthening collaboration between academia and industry is crucial to ensuring that SATD detection tools are not only scientifically rigorous but also practical and widely applicable in real-world software development settings.
+5.4 Decision-making guidance for adopting SATD detection techniques
+To address the lack of actionable guidance for adoption, this subsection synthesizes evidence from our review into practical decision considerations for selecting SATD detection techniques.
+Below, we summarize the findings based on the techniques, approaches, or models that were identified in our study.
+We also summarize the main advantages and disadvantages of each technique or approach.
+  • Heuristic-based techniques (e.g., tagging rules, regular expressions, pattern matching) are most appropriate when teams need a lightweight solution with minimal infrastructure and strong interpretability.
+They are typically easy to implement, explain, and integrate into existing workflows, making them suitable for small-to-medium codebases or for quick triage (e.g., flagging obvious TODO/FIXME-like debt markers).
+However, the review evidence indicates that heuristics often miss more subtle SATD expressions and are sensitive to vocabulary differences across projects and developer writing styles.
+Therefore, heuristics are best positioned as a baseline or as a high-precision filter when false positives are costly, rather than as a comprehensive detector.
+    - Advantages: Lightweight, easy to use
+    - Disadvantages: Lower recall, limited coverage
+  • Classic ML approaches (e.g., SVM, RF, and LR) offer a pragmatic middle ground for teams that have access to labeled data and seek improvements over heuristics without the full operational burden of deep models.
+Across the reviewed studies, ML methods tend to perform more robustly than heuristics while remaining easier to reproduce and debug than DL/Transformer approaches.
+ML is often a sensible choice for organizations that can maintain a modest labeling pipeline and require reasonably transparent features (e.g., TF-IDF-based models) to secure stakeholder buy-in.
+Nevertheless, reported performance remains highly dependent on dataset characteristics (e.g., class imbalance, artifact type, and preprocessing), and practitioners should expect re-tuning when moving to new domains.
+    - Advantages: More robust than heuristics, can be reasonably interpretable (feature-based)
+    - Disadvantages: Requires labeled data, performance may degrade under domain shift
+  • Deep learning approaches (e.g., CNN/LSTM/GNN) are most appropriate when the goal is improved effectiveness under complex linguistic variation or when large training corpora are available.
+The review evidence shows that DL methods often improve detection quality relative to ML in within-project evaluations, but they entail higher operational costs (training time, hardware, maintenance) and reduced transparency.
+DL models are therefore best suited to teams that can sustain the MLOps overhead, have sufficient data volume, and accept that model explanations may require additional techniques (e.g., post hoc explainability) to support developers’ trust.
+    - Advantages: Often higher effectiveness than classic ML, can leverage larger corpora and richer representations
+    - Disadvantages: Higher compute and MLOps cost, lower transparency
+  • Transformer-based models (e.g., BERT-family and related variants) are typically justified when SATD detection is a high-value, high-scale activity and teams prioritize effectiveness and broader language understanding.
+In the reviewed literature, Transformers often achieve the strongest reported results for SATD identification and are increasingly used across diverse textual artifacts.
+The main practical constraints are operational: they require more computational resources, more specialized expertise, and careful engineering for deployment and monitoring.
+In addition, their “black-box” nature can hinder adoption unless paired with interpretability and documentation practices.
+Practitioners should treat Transformers as a strategic investment: most suitable for large projects, multi-artifact pipelines, or scenarios in which missing SATD instances (false negatives) incur high downstream costs.
+    - Advantages: State-of-the-art effectiveness, strong language understanding across artifacts
+    - Disadvantages: Highest compute and expertise requirements, interpretability and deployment complexity
+These qualitative observations are consistent with the overall performance patterns synthesized in Section 4.3, where more recent DL and Transformer-based methods tend to report higher precision/recall/F1 than earlier heuristic and classical ML approaches, albeit under heterogeneous datasets and evaluation protocols.
+Our synthesis indicates that no single technique dominates across all criteria; the best choice depends on project constraints (e.g., available labeled data, compute budget, and expertise) and the intended use of the detector (e.g., quick triage vs.
+high recall detection).
+Based on the cross-study synthesis, practitioners can use the following cues to narrow down their choice: if the goal is immediate, low-cost triage with maximum explainability, heuristic-based techniques are typically the best starting point.
+If labeled data are available and reproducibility and maintainability are important, classic ML provides a strong and practical baseline.
+If performance requirements are higher and sufficient data/compute resources exist, DL techniques become appropriate.
+Finally, if an organization can support the operational overhead and needs state-of-the-art effectiveness across heterogeneous artifacts, Transformer-based models are generally the most suitable option.
+5.5 Implications for practitioners
+BERT-based models demonstrate superior precision and recall in SATD detection: by adopting these models, practitioners can significantly reduce their reliance on manual review processes, which are often time-consuming, inconsistent, and prone to human error.
+Automated approaches streamline the identification of technical debt, ensuring that critical SATD instances are accurately detected without the extensive overhead of manual analysis.
+This enables software teams to allocate their resources more efficiently and to focus on higher-value activities, such as refactoring and feature development.
+However, one of the key challenges in SATD identification is class imbalance, where SATD instances are significantly outnumbered by non-SATD instances.
+This imbalance often leads detection models to favor the dominant (non-SATD) class, resulting in missed detections.
+Additionally, class imbalance affects SATD categorization, as SATD must be classified into specific types of technical debt, such as design debt, architecture debt, requirements debt, and others.
+In both cases, the imbalance increases the risk of false negatives 117, where important SATD instances are overlooked.
+To mitigate this issue, practitioners should employ strategies such as data augmentation (e.g., by creating additional synthetic SATD samples to balance the dataset) and domain-adaptation techniques, which optimize model performance across diverse projects and environments.
+These approaches help ensure that models maintain high accuracy and recall, even when faced with the imbalanced data typical of real-world software projects.
+As outlined in subsection 5.4, we present a comprehensive framework to assist practitioners in selecting the most appropriate SATD detection approach tailored to their specific project needs and constraints.
+This framework enables practitioners to effectively balance trade-offs involving scalability, explainability, and resource demands.
+These factors are essential for the successful adoption of SATD detection tools.
+Additionally, automated SATD detection and early identification of technical debt contribute to the sustainability of software systems.
+Since SATD is prevalent across many software projects (affecting between 2.4% and 31% of files 8), it presents a long-term challenge.
+SATD often persists for extended periods, with a median lifespan ranging from 18 to 172 days and, in some cases, lasting through more than 1,000 commits 123.
+By adopting automated detection techniques, organizations that adopt these techniques will likely see measurable improvements in development efficiency, reduced technical debt accumulation, and enhanced product quality.
+These benefits help mitigate the risks associated with delayed debt repayment.
 5.6 Future directions
@@
 At the same time, incorporating explainable AI (XAI) techniques is crucial to ensure that the decision-making processes of these models are transparent and understandable for developers.
-Developing more interpretable models, or employing post-hoc explanation techniques, could foster trust in these tools and encourage their adoption in real-world projects.
-Further research on enhancing the interpretability of deep learning and Transformer models—potentially through visualizations or simplified intermediate representations—would empower practitioners to better understand, trust, and effectively utilize these models in software development settings.
-There is also a need for the development of more practical SATD detection tools that can be easily integrated into existing development environments.
-These tools should focus on providing real-time feedback to developers, enabling them to address technical debt as part of their normal workflow.
-By bridging the gap between research and practice, the software engineering community can ensure that SATD detection techniques contribute meaningfully to the long-term maintenance and evolution of software systems.
+Developing more interpretable models or employing post-hoc explanation techniques could foster trust in these tools and encourage their adoption in real-world projects.
+Future work should focus on improving the interpretability of deep learning and Transformer models, for example, by using visualizations, highlighting informative patterns, and providing simplified intermediate representations.
+These improvements would help practitioners better understand the models, build trust in their outputs, and use them more effectively in software development contexts.
+Beyond improving models, future research should prioritize SATD detection tools that are scalable, easy to use, and designed for real-world adoption.
+In practice, this means tools that integrate smoothly with existing development environments, support team collaboration, and work across diverse software ecosystems, including multi-language codebases.
+Since modern workflows are centered on IDEs and CI/CD pipelines, SATD detectors should be deployable in both settings and provide timely, actionable signals.
+In particular, real-time or near-real-time feedback can help developers identify and address technical debt as part of their normal workflow, rather than as a follow-up activity performed after implementation.
+By closing the gap between research prototypes and production-ready tooling, the software engineering community can ensure that SATD detection techniques deliver tangible value for long-term software maintenance and evolution.
+While our findings provide valuable insights into SATD detection, it is essential to acknowledge potential limitations that may influence the generalizability of our conclusions.
+The next section discusses threats to validity, including biases in study selection, dataset heterogeneity, and variations in evaluation metrics, and how we have addressed these concerns.
 6 Threats to Validity
-As with any systematic literature review (SLR), the results of this study are subject to various validity threats.
+As with any systematic literature review, the results of this study are subject to various validity threats.
 These are categorized as internal, external, and construct validity threats, along with mitigation strategies to minimize their impact.
@@
 One key threat to internal validity in this SLR is related to the formulation of the search strings.
-In the pilot search, several iterations of search strings were tested, starting with a more general search (e.g., “self-admitted technical debt" OR SATD and NLP), which led to a broader, less focused set of results.
-Over time, more refined strings were developed, including terms like “detection" and “identification," as well as the use of boolean operators and the wildcard (*) for broader term variations.
-The final search string (S3), which combined terms such as “self-admitted technical debt" OR SATD with NLP OR “natural language processing" AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies.
+In the pilot search, several iterations of search strings were tested, starting with a more general query (e.g., “self-admitted technical debt” OR SATD and NLP), which yielded a broader, less focused set of results.
+Over time, more refined strings were developed, including terms such as “detection” and “identification”, as well as the use of Boolean operators and the wildcard (*) for broader term variations.
+The final search string (S3), which combined terms such as “self-admitted technical debt” OR SATD with NLP OR “natural language processing” AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies.
 This iterative process helped mitigate the risk of missing key studies during the initial literature search.
@@
 A primary threat to external validity in this research is the limited generalizability of the results due to the scope of the reviewed datasets.
-All of the studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints faced in proprietary or large-scale industrial environments.
+All studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints encountered in proprietary or large-scale industrial environments.
 Open-source datasets often have characteristics, such as well-documented codebases or community-driven development practices, that differ significantly from proprietary software projects, which may involve stricter deadlines, unique domain requirements, or higher security and compliance constraints.
 As a result, some of the trends and findings related to SATD detection may not extend to these more complex and regulated environments.
-For example, the performance of machine learning or deep learning models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
+For example, the performance of ML or DL models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
+Another potential limitation of this review is that the synthesized results for RQ3 are based exclusively on within-project evaluations (e.g., k-fold cross-validation or random splits on the same project).
+While this choice ensures comparability across studies, it may overestimate performance relative to cross-project or out-of-distribution evaluations, which better reflect real-world deployment scenarios.
+Some primary studies did report cross-project results, but these were not included in our synthesis.
+Therefore, the performance metrics presented in this review should be interpreted with caution, as they represent results under controlled within-project settings rather than true generalizability across projects.
 6.3 Construct validity
-Construct validity pertains to whether the study accurately measures the concepts it aims to investigate.
-A significant threat in this SLR is the comparison of results from different studies, which can be misleading if the studies use different methodologies or report metrics inconsistently.
-Directly comparing the precision, recall, and F1-scores of models trained on heterogeneous datasets without considering their context could lead to inaccurate conclusion.
-To address this, we employed a grouping strategy to compare similar studies and ensured that the datasets were organized before any comparisons were made.
+Construct validity concerns whether a study accurately measures the concepts it aims to investigate.
+A significant threat in this SLR is the comparison of results across studies, which can be misleading if studies employ different methodologies or report metrics inconsistently.
+Directly comparing the precision, recall, and F1 scores of models trained on heterogeneous datasets without considering their context may lead to inaccurate conclusions.
+To address this, we employed a grouping strategy to compare similar studies and ensured that the SATD tasks were organized before any comparisons were made.
 By contextualizing the results within groups, we aimed to make the comparisons more meaningful.
-Furthermore, this methodology is based on established practices within the field, notably exemplified by Hall et al.
-103 and Malhotra 104, which similarly undertakes the task of comparing results across different studies to deduce broad-based insights into algorithmic performance in varying contexts.
+Furthermore, this methodology is based on established practices within the field, exemplified by Hall et al.
+112 and Malhotra 113, which similarly compare results across studies to derive broad-based insights into algorithmic performance across varying contexts.
+Having considered the potential limitations of this study, we now summarize our key findings and contributions.
+The final section revisits the research questions, highlights the main takeaways, and outlines directions for future work to advance SATD detection research and tool development.
 7 Conclusion
-This SLR provides an in-depth analysis of SATD detection approaches over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated machine learning and deep learning models.
-Among the key findings, it is evident that while NLP-based techniques were foundational in early SATD detection, they struggle with recall and overall performance in comparison to ML and DL approaches.
-Transformer-based models, particularly BERT, emerged as the most effective in capturing complex patterns in textual artifacts, offering superior performance across multiple metrics.
-However, several challenges persist, including issues related to dataset diversity, the scalability of models, and the explainability of AI techniques.
-The reliance on open-source datasets in many studies limits the generalizability of the findings, and more research is needed to adapt these models to varied industrial environments.
-Moreover, the imbalance of SATD datasets continues to affect the robustness of model performance, and integrating multiple textual artifacts such as commit messages and pull requests presents further complexity.
-Furthermore, the lack of widely adopted tools for SATD detection in industry highlights a gap between research advancements and practical applications.
-In conclusion, while the field of SATD detection has advanced considerably, there remains a need for further research and collaboration between academia and industry to ensure these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
+This SLR provides an in-depth analysis of SATD detection techniques over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated ML, DL, and Transformer-based models.
+Among the key findings, it is evident that while heuristic-based techniques were foundational in early SATD detection, they struggle to achieve comparable recall and overall performance to ML and DL techniques.
+Transformer-based models, particularly BERT, have emerged as the most effective for capturing complex patterns in textual artifacts, achieving superior performance across multiple metrics.
+However, several challenges persist, including computational costs, model scalability, and the explainability of AI techniques.
+The reliance on open-source datasets in many studies limits the generalizability of findings, and further research is needed to adapt these models to diverse industrial environments.
+Moreover, the imbalance in SATD datasets continues to affect model robustness, and integrating multiple textual artifacts, such as commit messages and pull requests, introduces further complexity.
+Furthermore, the lack of widely adopted tools for SATD detection in the industry highlights a gap between research advancements and practical applications.
+In conclusion, while the field of SATD detection has advanced considerably, further research and collaboration between academia and industry are needed to ensure that these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
 […]
-[…]
-[…]
+*ORCID Edi Sutoyo https://orcid.org/0000-0002-8413-5070
+Andrea Capiluppi https://orcid.org/0000-0001-9469-6050
 […]
@@
 […]
-*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Financing of Higher Education (BPPT) and Indonesia Endowment Fund for Education (LPDP).
+*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Higher Education Funding and Assessment (PPAPT), Indonesia Endowment Fund for Education (LPDP), and the Ministry of Higher Education, Science, and Technology of the Republic of Indonesia.